### v2

# Dynamic Programming: From Local Optimality to Global Optimality

Abstract.
In the theory of dynamic programming, an optimal policy is a policy whose lifetime value dominates that of all other policies from every possible initial condition in the state space. This raises a natural question: when does optimality from a single state imply optimality from every state? We show that, in a general setting, irreducibility of the transition kernel is sufficient for this property. Our results have important implications for modern policy-based algorithms used to solve large-scale dynamic programs in reinforcement learning and other fields.
[…]
1. Introduction
Dynamic programming (DP) is a major branch of optimization theory, with applications ranging from fine tuning of large language models to DNA sequencing, space exploration, and air traffic control. Dynamic programs that include uncertainty are often called Markov decision processes (MDPs) and the theory of such processes has been extensively developed (see, e.g., Bäuerle and Rieder (2011), Hernández-Lerma and Lasserre (2012), Bertsekas (2012), or Bertsekas (2022)). Much of the recent surge in interest in MDPs has been fueled by artificial intelligence and reinforcement learning (see, e.g., Bertsekas (2021) or Kochenderfer et al. (2022)).
In recent years, researchers solving large-scale MDPs have moved away from value-based methods and towards policy-based methods, one example of which is policy gradient ascent (see, e.g., Murphy (2024), Sutton et al. (1999), Lan et al. (2023), Kumar et al. (2023), or Friedl et al. (2023)). These policy-based methods seek to maximize a real-valued objective, such as
$m(\sigma)\coloneq\int v_{\sigma}(x)\rho(\mathop{}\!\mathrm{d}x)\qquad(\sigma\in\Sigma).$ (1)
Here $\sigma$ is a policy for a given MDP, mapping some state space $\mathsf{X}$ into a specified action space, $v_{\sigma}(x)$ represents the lifetime value of following the fixed policy $\sigma$, conditional on initial state $x$, and $\rho$ is a given “initial distribution.” In practice, each $\sigma$ is typically represented by a neural network. Policy-based methods often handle large problems with continuous state and action spaces more efficiently than traditional value-based methods such as value function iteration (VFI). The framework has led to numerous successful algorithms for solving complex decision-making problems, including Trust Region Policy Optimization (Schulman et al., 2015), Asynchronous Advantage Actor-Critic (Mnih et al., 2016), and Proximal Policy Optimization (Schulman et al., 2017).
Despite these successes, there is one significant disadvantage of policy gradient methods: unlike some traditional DP algorithms, these methods are not guaranteed to find an optimal policy. To understand the issues at hand, recall that an optimal policy is a feasible policy $\sigma$ such that $v_{\sigma}(x)=\max_{s\in\Sigma}v_{s}(x)$ for every $x\in\mathsf{X}$; that is, a policy $\sigma$ such that following this policy in every period leads to maximum lifetime value from every initial state $x$. Even if one attains a global maximum in (1), with maximization over all $\sigma\in\Sigma$, there is no guarantee that the resulting policy will be an optimal policy.
The prevailing view is that obtaining an approximately optimal policy depends heavily on the initial distribution $\rho$ that is used to construct the objective function (1). For example, in an important study of policy gradient methods, Bhandari and Russo (2024) state that “Policy gradient methods have poor convergence properties if applied without an exploratory initial distribution” (Bhandari and Russo, 2024, p. 1910). Here “exploratory” means that $\rho$ in (1) should have a large support.
One disadvantage of choosing $\rho$ with large support is that this exploration can be computationally expensive. This motivates the following question: When is it permissible to choose $\rho$ with small support? At the extreme, when can $\rho$ be concentrated at a single point $x$, so that the maximization criterion $m(\sigma)$ from (1) is just $v_{\sigma}(x)$, and yet maximization of this criterion over all $\sigma\in\Sigma$ yields a (globally) optimal policy? In other words, when does optimality at a single state imply optimality at every state? We show that, for standard MDPs on general state spaces, irreducibility of the Markov dynamics generated by $\sigma$ is sufficient for this property. Specifically, if a policy is optimal at a single state and has an irreducible transition kernel, then this optimality propagates throughout the entire state space, making the policy globally optimal. Similarly, if irreducibility holds and there exists a distribution $\rho$ such that $\int v_{\sigma}\mathop{}\!\mathrm{d}\rho=\max_{s\in\Sigma}\int v_{s}\mathop{}\!\mathrm{d}\rho$, then $\sigma$ is an optimal policy. In addition, we show that two weak forms of irreducibility are sufficient for this result when $\sigma$ is also continuous.
Since gradient policy methods are commonly applied to problems with continuous state and action spaces, we avoid discreteness restrictions. In particular, for the MDPs that we consider, the state and action spaces can be arbitrary metric spaces. We also avoid placing restrictions on the class of MDPs under consideration, adopting only mild regularity conditions that imply existence of solutions. Focusing on problems where solutions exist is natural for this line of research, since we wish to examine conditions under which local optima imply global optima.
Papers examining theoretical properties of policy gradient methods have some connection to our work. For example Bhandari and Russo (2024) examine when policy gradient ascent (actually decent) yields a globally optimal policy. One the one hand, Bhandari and Russo (2024) directly examine the policy gradient algorithm and find sharp new results for several important cases. On the other hand, to step from local to global optimality, they restrict the classes of MDPs under consideration and assume a large support for $\rho$. Here we avoid large support restrictions on $\rho$ (since imposing such restrictions directly imposes a connection from local to global optimality—see Lemma 3.2 and the surrounding discussion).
Other papers that examine theoretical properties of gradient policy methods include Khodadadian et al. (2021), Agarwal et al. (2021), and Xiao (2022). However, in these papers, the focus is on proving the convergence of $\int v_{\sigma}\mathop{}\!\mathrm{d}\rho$ to its maximal value, rather than obtaining global convergence from local convergence (as we do here). At the same time, these papers provide valuable rates of convergence for specific algorithms, which we do not discuss. A related line of research focuses on average-optimal policies in finite-state MDPs by leveraging specific state space structures under some policies. This includes exploring unichain, multichain, communicating, and weakly communicating MDPs to study algorithmic convergence (Bartlett and Tewari, 2009; Puterman, 2014). Our results also demonstrate that, in finite-state MDPs, optimality can extend from a single state to all accessible states. Thus, our results are applicable to various classes of MDPs in this line of research and support the development of more efficient algorithms.
The paper is structured as follows. Section 2 provides background on MDPs. Section 3 states our main results in a general setting. Section 4 examines how the irreducibility condition from Section 3 can be weakened while still obtaining some transmission of optimality across states. Section 5 illustrates our theoretical results in the context of a benchmark optimal savings problem. Section 6 outlines avenues for future work.
2. Markov Decision Processes
In this section, we review essential properties of Markov decision processes and state a technical lemma that will be applied in our main results.
2.1. Preliminaries
[…]
A transition kernel on $\mathsf{X}$ is a function $P$ from $\mathsf{X}\times\mathscr{B}$ to $[0,1]$ such that $x\mapsto P(x,B)$ is Borel measurable for all $B\in\mathscr{B}$ and $B\mapsto P(x,B)$ is a Borel probability measure for all $x\in\mathsf{X}$. To any such transition kernel $P$ we associate a bounded linear operator on $b\mathsf{X}$, often referred to as its Markov operator and also denoted by $P$, via
$f\mapsto Pf,\qquad(Pf)(x)=\int f(x^{\prime})P(x,\mathop{}\!\mathrm{d}x^{\prime}).$ (2)
In what follows, $(Pf)(x)$ will be understood as the expectation of $f(X_{t+1})$ given that $X_{t}=x$ and $X_{t+1}$ is drawn from $P(x,\mathop{}\!\mathrm{d}x^{\prime})$.
As usual, the positive cone of $b\mathsf{X}$, denoted here by $b\mathsf{X}_{+}$, is the set of all nonnegative functions in $v\in b\mathsf{X}$. Let $b\mathsf{X}^{\prime}$ be the dual space of $b\mathsf{X}$ and let $b\mathsf{X}^{\prime}_{+}$ to be the positive cone of $b\mathsf{X}^{\prime}$. The set $b\mathsf{X}^{\prime}_{+}$ contains, among other objects, the set $\mathcal{D}(\mathsf{X})$ of Borel probability measures on $\mathsf{X}$. For simplicity, elements of $\mathcal{D}(\mathsf{X})$ are referred to as distributions. For $\rho\in\mathcal{D}(\mathsf{X})$ and $f\in b\mathsf{X}$ we set
$\left\langle{f,\rho}\right\rangle\coloneq\int f\mathop{}\!\mathrm{d}\rho.$
For each $x\in\mathsf{X}$, the point evaluation functional generated by $x$ is the map $\delta_{x}$ that sends each $w\in b\mathsf{X}$ into $w(x)\in\mathbbm{R}$. Below it will be convenient for us to write this in dual notation, so that $\left\langle{w,\delta_{x}}\right\rangle=w(x)$ for every $w\in b\mathsf{X}$. We will make use of the following lemma.
Lemma 2.1.
Every point evaluation functional on $b\mathsf{X}$ is a nonzero element of $b\mathsf{X}_{+}^{\prime}$.
Proof.
Fix $x\in\mathsf{X}$. Linearity of $\delta_{x}$ is obvious: given $a,b\in\mathbbm{R}$ and $v,w\in b\mathsf{X}$, we have
$\left\langle{av+bw,\delta_{x}}\right\rangle=(av+bw)(x)=av(x)+bw(x)=a\left\langle{v,\delta_{x}}\right\rangle+b\left\langle{w,\delta_{x}}\right\rangle.$
Regarding continuity, if $w_{n}\to w$ in $b\mathsf{X}$, then $w_{n}\to w$ pointwise on $\mathsf{X}$, so $\left\langle{w_{n},\delta_{x}}\right\rangle=w_{n}(x)\to w(x)=\left\langle{w,\delta_{x}}\right\rangle$. Regarding positivity, it suffices to show that $\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$ whenever $w\geqslant 0$. This clearly holds, since $w\geqslant 0$ implies $w(x)=\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$. Finally, $\delta_{x}$ is not the zero element of $b\mathsf{X}^{\prime}$ because we can always take a $w=\mathbf{1}\in b\mathsf{X}$ with $\left\langle{w,\delta_{x}}\right\rangle=w(x)=1\not=0$. ∎
2.2. Markov Decision Process
Let $\mathsf{X}$ and $\mathsf{A}$ be metric spaces, as in Section 2.1. A Markov decision process (MDP) with state space $\mathsf{X}$ and action space $\mathsf{A}$ is a tuple $(r,\Gamma,\beta,P)$, where $r$ is a reward function, $x\mapsto\Gamma(x)\subset\mathsf{A}$ is a feasible correspondence, $\beta$ is a discount factor and $P(x,a,\mathop{}\!\mathrm{d}x^{\prime})$ is a distribution over next period states given current state $x$ and action $a$. Let $\mathsf{G}\coloneq\{(x,a)\in\mathsf{X}\times\mathsf{A}:a\in\Gamma(x)\}$. We consider a relatively standard environment, as considered in, say, Bäuerle and Rieder (2011) and Hernández-Lerma and Lasserre (2012), where
  1. (a)
    $\beta\in(0,1)$,
  2. (b)
    $\Gamma$ is nonempty, continuous, and compact-valued on $\mathsf{X}$,
  3. (c)
    $r$ is bounded and continuous on $\mathsf{G}$, and
  4. (d)
    the map $(x,a)\mapsto\int v(x^{\prime})P(x,a,dx^{\prime})$ is continuous on $\mathsf{G}$ whenever $v\in bc\mathsf{X}$.
(The case of unbounded $r$ is discussed in Section 6.)
Let $\Sigma$ denote the set of feasible policies, by which we mean all Borel measurable functions $\sigma$ mapping $\mathsf{X}$ to $\mathsf{A}$ with $\sigma(x)\in\Gamma(x)$ for all $x\in\mathsf{X}$. For each $\sigma\in\Sigma$ and $x\in\mathsf{X}$, we set
$r_{\sigma}(x)\coloneq r(x,\sigma(x))\quad\text{and}\quad P_{\sigma}(x,\mathop{}\!\mathrm{d}x^{\prime})\coloneq P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime}).$
Thus, $r_{\sigma}(x)$ is rewards at $x$ under policy $\sigma$ and $P_{\sigma}$ is the transition kernel on $\mathsf{X}$ generated by $\sigma$. Using the corresponding Markov operator $P_{\sigma}$, as defined in (2), the lifetime value of a policy $\sigma$, denoted by $v_{\sigma}$, can be expressed as
$v_{\sigma}=\sum_{t=0}^{\infty}(\beta P_{\sigma})^{t}r_{\sigma}=(I-\beta P_{\sigma})^{-1}r_{\sigma}.$ (3)
(See, e.g., Puterman (2014), Theorem 6.1.1.) The lifetime value $v_{\sigma}$ defined in (3) is the unique fixed point in $b\mathsf{X}$ of the policy operator $T_{\sigma}$ defined by $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$. This operator can be written more explicitly as
$(T_{\sigma}\,v)(x)=r(x,\sigma(x))+\beta\int v(x^{\prime})P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime})\qquad(v\in b\mathsf{X},\;x\in\mathsf{X}).$
(The lifetime value $v_{\sigma}$ is the unique fixed point of $T_{\sigma}$ because the spectral radius of the linear operator $\beta P_{\sigma}$ is $\beta$, so $\beta<1$ implies that $v=r_{\sigma}+\beta P_{\sigma}\,v$ has the unique solution given by the right-hand side of (3).) Iterating on the definition $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$, we find that
$T_{\sigma}^{n}v=r_{\sigma}+\beta P_{\sigma}r_{\sigma}+\cdots+(\beta P_{\sigma})^{n-1}r_{\sigma}+(\beta P_{\sigma})^{n}v\quad\text{for all $n\in\mathbbm{N}$}.$ (4)
This expression will be useful in the theory below.
The value function is denoted $v^{*}$ and defined at each $x\in\mathsf{X}$ by $v^{*}(x)\coloneq\sup_{\sigma\in\Sigma}v_{\sigma}(x)$. A policy $\sigma$ is called optimal if $v_{\sigma}(x)=v^{*}(x)$ for all $x\in\mathsf{X}$.
We define the Bellman operator by
$(Tv)(x)=\max_{a\in\Gamma(x)}\left\{r(x,a)+\beta\int v(x^{\prime})P(x,a,\mathop{}\!\mathrm{d}x^{\prime})\right\}\qquad(v\in b\mathsf{X},\;x\in\mathsf{X}).$ (5)
We will use the following facts:
Proposition 2.2.
Under the stated assumptions,
  1. (a)
    the value function $v^{*}$ is the unique fixed point of the Bellman operator in $b\mathsf{X}$
  2. (b)
    the value function $v^{*}$ is well-defined and contained in $bc\mathsf{X}$ and
  3. (c)
    at least one optimal policy exists.
Proof.
See Hernández-Lerma and Lasserre (2012) or Bäuerle and Rieder (2011). ∎
We also make use of the following technical lemma, which shows one implication of local optimality at some given $x\in\mathsf{X}$.
Lemma 2.3.
If $\sigma\in\Sigma$ and $v_{\sigma}(x)=v^{*}(x)$, then
$\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})=0\qquad\text{for all }n\in\mathbbm{N}.$ (6)
Proof.
Fix $n\in\mathbbm{N}$. Applying the expression for $T_{\sigma}^{n}w$ from (4) twice, first with $v=v_{\sigma}$ and then with $v=v^{*}$, we get
$T_{\sigma}^{n}\,v_{\sigma}-T_{\sigma}^{n}\,v^{*}=\beta^{n}(P_{\sigma}^{n}\,v_{\sigma}-P_{\sigma}^{n}\,v^{*}).$ (7)
In addition, we have
$v_{\sigma}=T_{\sigma}^{n}\,v_{\sigma}\leqslant T_{\sigma}^{n}\,v^{*}\leqslant T^{n}\,v^{*}=v^{*}.$ (8)
In (8), the first inequality is due to the fact $T_{\sigma}$ is order preserving and $v_{\sigma}\leqslant v^{*}$, while the second follows from the fact that $T_{\sigma}\,v\leqslant Tv$ for all $v\in b\mathsf{X}$. (Since $T_{\sigma}\,v\leqslant T\,v$ for all $v$, $T_{\sigma}^{2}\,v^{*}\leqslant T_{\sigma}T\,v^{*}\leqslant T^{2}\,v^{*}$ and so $T_{\sigma}^{2}\,v^{*}\leqslant T^{2}\,v^{*}.$ By induction, the second inequality holds.) The claim in Lemma 2.3 follows from (7) and (8). To see this, fix $x\in\mathsf{X}$ with $v_{\sigma}(x)=v^{*}(x)$. From this equality and (8) we get $(T_{\sigma}^{n}\,v_{\sigma})(x)=(T_{\sigma}^{n}\,v^{*})(x)$. Since $\beta>0$, combining this result with (7) yields $(P_{\sigma}^{n}\,v_{\sigma})(x)=(P_{\sigma}^{n}\,v^{*})(x)$. Hence (6) holds. ∎
3. From Local to Global Optimality
The standard definition of optimality, which was given in Section 2.2, is global in nature, since it concerns the lifetime value of the policy at every $x\in\mathsf{X}$. We seek conditions under which local optimality implies global optimality. In particular, we seek conditions under which the following three statements are equivalent:
  1. (E1)
    there exists an $x\in\mathsf{X}$ such that $v_{s}(x)\leqslant v_{\sigma}(x)$ for all $s\in\Sigma$,
  2. (E2)
    there exists a $\rho\in\mathcal{D}(\mathsf{X})$ such that $\left\langle{v_{s},\rho}\right\rangle\leqslant\left\langle{v_{\sigma},\rho}\right\rangle$ for all $s\in\Sigma$,
  3. (E3)
    $\sigma$ is an optimal policy.
3.1. Preliminary Results
We first note that, in the MDP set up we have described, (E1) and (E2) are always equivalent, as the next lemma shows.
Lemma 3.1.
If $\sigma$ is a feasible policy, then the statements (E1) and (E2) are equivalent.
Proof.
To show (E1) implies (E2), assume (E1) and fix $x\in\mathsf{X}$ with $v_{\sigma}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$. Let $\delta_{x}$ be the point evaluation functional generated by $x$. Since $\delta_{x}\in\mathcal{D}(\mathsf{X})$ and $\left\langle{v_{\sigma},\delta_{x}}\right\rangle=v_{\sigma}(x)\geqslant v_{s}(x)=\left\langle{v_{s},\delta_{x}}\right\rangle$ for all $s\in\Sigma$, so (E2) holds. To show (E2) implies (E1), fix $\rho\in\mathcal{D}(\mathsf{X})$ with $\left\langle{v_{\sigma},\rho}\right\rangle\geqslant\left\langle{v_{s},\rho}\right\rangle$ for all $s\in\Sigma$. Suppose to the contrary that for each $x\in\mathsf{X},$ we can find a $\tau\in\Sigma$ such that $v_{\tau}(x)>v_{\sigma}(x)$. Since $v^{*}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$ and $x\in\mathsf{X}$, we have $v^{*}(x)>v_{\sigma}(x)$ for all $x\in\mathsf{X}$. Hence $\left\langle{v^{*},\rho}\right\rangle>\left\langle{v_{\sigma},\rho}\right\rangle$. This contradiction proves (E1). ∎
Our second preliminary observation suggests one way to obtain (E3) from (E2).
Lemma 3.2.
If $\left\langle{v_{\sigma},\rho}\right\rangle=\left\langle{v^{*},\rho}\right\rangle$, then $v_{\sigma}=v^{*}$ holds $\rho$-almost everywhere.
Proof.
By definition, $v^{*}\geqslant v_{\sigma}$ on $\mathsf{X}$. If, in addition, $v^{*}>v_{\sigma}$ on a set $E$ of positive $\rho$-measure, then
$\int(v^{*}-v_{\sigma})\mathop{}\!\mathrm{d}\rho\geqslant\int_{E}(v^{*}-v_{\sigma})\mathop{}\!\mathrm{d}\rho>0.$
This contradicts $\left\langle{v_{\sigma},\rho}\right\rangle=\left\langle{v^{*},\rho}\right\rangle$. Hence $v_{\sigma}=v^{*}$ holds $\rho$-almost everywhere. ∎
To illustrate Lemma 3.2, suppose that $\mathsf{X}$ is discrete. In this case, the lemma tells us that $\left\langle{v_{\sigma},\rho}\right\rangle=\left\langle{v^{*},\rho}\right\rangle$ implies $v_{\sigma}=v^{*}$ (i.e., (E3) holds) when $\rho$ is supported on all of $\mathsf{X}$. In other words, maximizing the scalar performance measure $\left\langle{v_{\sigma},\rho}\right\rangle$ with a highly exploratory initial distribution $\rho$ guarantees global optimality.
The disadvantage of this approach is that evaluating the scalar measure from a highly exploratory initial distribution is costly. For this reason, our main interest is in providing conditions under which (E1) implies (E3). Our conditions relate to irreducibility. The first condition, discussed in Section 3.2, uses a notion of irreducibility from the literature on Banach lattices (see, e.g., Zaanen (2012)), while the second, discussed in Section 4.3, uses an irreducibility concept from the Markov process literature.
3.2. Strong Irreducibility
Recall that a linear subspace $I$ of $b\mathsf{X}$ is called an ideal in $b\mathsf{X}$ when $f\in I$ and $|g|\leqslant|f|$ implies $g\in I$. An ideal $I$ is said to be invariant for a linear operator $K$ if $KI\subset I$. A linear operator $K$ from $b\mathsf{X}$ to itself is called positive when $Kf\geqslant 0$ for all $f\geqslant 0$. A positive linear operator $K$ is called irreducible if the only invariant ideals under $K$ are the trivial subspace $\{0\}$ and the whole space $b\mathsf{X}$. We will make use of the characterization in Proposition 8.3 (c) of Schaefer (1974): A positive linear operator $K$ on $b\mathsf{X}$ is irreducible if and only if, for each nonzero $f\in b\mathsf{X}_{+}$ and each nonzero $\mu\in b\mathsf{X}^{\prime}_{+}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mu,K^{m}f}\right\rangle>0$.
In what follows, we call a transition kernel $P$ on $\mathsf{X}$ strongly irreducible if its Markov operator (see (2)) is irreducible on $b\mathsf{X}$ in the sense just defined.
Here is our main result for the strongly irreducible case. In the statement $\sigma$ is any feasible policy.
Theorem 3.3.
If $P_{\sigma}$ is strongly irreducible, then (E1)(E3) are equivalent.
For example, Theorem 3.3 tells us that, under the stated conditions, we can obtain an optimal policy by fixing an arbitrary initial state $x\in\mathsf{X}$ and maximizing the real-valued function $s\mapsto v_{s}(x)$ over $\Sigma$. Alternatively, we can fix any distribution $\rho$ and maximize $s\mapsto\left\langle{v_{s},\rho}\right\rangle$.
Proof of Theorem 3.3.
In view of Lemma 3.1, it suffices to show that (E1) and (E3) are equivalent. That (E3) implies (E1) is immediate from the definition of optimal policies. Hence we need only show that (E1) implies (E3). In line with the conditions of Theorem 3.3, we assume that $\sigma$ is a feasible policy and $P_{\sigma}$ is strongly irreducible.
Let $h\coloneq v^{*}-v_{\sigma}$. By the definition of $v^{*}$ we have $0\leqslant h$. We claim in addition that $h=0$. To see this, suppose to the contrary that $h$ is nonzero. In this case, by strong irreducibility, for each nonzero $\mu$ in the positive cone of $b\mathsf{X}^{\prime}$ we can find an $m\in\mathbbm{N}$ such that $\left\langle{\mu,P_{\sigma}^{m}\,h}\right\rangle>0$. Because $\delta_{\bar{x}}$ is a nonzero element of the positive cone of $b\mathsf{X}^{\prime}$, we can set $\mu=\delta_{\bar{x}}$ to obtain an $m\in\mathbbm{N}$ with $(P_{\sigma}^{m}\,h)(\bar{x})>0$. This contradicts (6), so $h=0$ holds. In other words, $v_{\sigma}(x)=v^{*}(x)$ for all $x\in\mathsf{X}$, as was to be shown. ∎
3.3. Finite States and Actions
In this section, we specialize to the case where $\mathsf{X}$ is finite and study the role of irreducibility in this setting. Note that the previous results (in particular, Theorem 3.3), can be applied by taking the metric on $\mathsf{X}$ to be the discrete metric, so that all subsets of $\mathsf{X}$ are open and $\mathscr{B}$ is the set of all subsets of $\mathsf{X}$. We also assume that $\mathsf{A}$ is finite and impose the discrete metric on $\mathsf{A}$.
Given a transition kernel $Q$ on $\mathsf{X}$, we say that $y$ in $\mathsf{X}$ is $Q$-accessible from $x\in\mathsf{X}$ when there exists an $m\in\mathbbm{N}$ such that $Q^{m}(x,y)>0$. The usual definition of irreducibility of a transition kernel $Q$ on finite $\mathsf{X}$ is that, for every $x,y\in\mathsf{X}$, $x$ is $Q$-accessible from $y$ and $y$ is $Q$-accessible from $x$. To distinguish between different notions of irreducibility, we call this discrete irreducibility.
Lemma 3.4.
Let $\sigma$ be any feasible policy. When the state space is finite, the following statements are equivalent:
  1. (a)
    $P_{\sigma}$ is strongly irreducible.
  2. (b)
    $P_{\sigma}$ is discretely irreducible.
Proof.
((a) $\implies$ (b)) Fix $x,y\in\mathsf{X}$. Let $\mathbbm{1}_{x}(z)$ equal $1$ when $z=0$ and zero elsewhere. Let $\mathbbm{1}_{y}$ be defined analogously. Since $\mathbbm{1}_{x}\in bm\mathsf{X}_{+}$ and $\mathbbm{1}_{y}\in(bm\mathsf{X}_{+})^{\prime}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle>0$. But $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle=P_{\sigma}^{m}(x,y)$, so $y$ is $P_{\sigma}$-accessible from $x$. This proves that $P_{\sigma}$ is discretely irreducible.
((b) $\implies$ (a)) Fix a nonzero $f\in bm\mathsf{X}_{+}$ and nonzero $\mu\in(bm\mathsf{X}_{+})^{\prime}$. Since $\mathsf{X}$ is finite, these element are just maps from $\mathsf{X}$ to $\mathbbm{R}_{+}$ and, as both are nonzero, we can find $\bar{x},\bar{y}\in\mathsf{X}$ such that $f(\bar{y})>0$ and $\mu(\bar{x})>0$. Moreover, since $P_{\sigma}$ is discretely irreducible, there is $m\in\mathbbm{N}$, such that $P_{\sigma}^{m}(\bar{x},\bar{y})>0$. As a result,
$\left\langle{\mu,P_{\sigma}^{m}f}\right\rangle=\sum_{x}(P_{\sigma}^{m}f)(x)\mu(x)=\sum_{x}\sum_{y}f(y)P_{\sigma}^{m}(x,y)\mu(x)\geqslant P_{\sigma}^{m}(\bar{x},\bar{y})f(\bar{y})\mu(\bar{x})>0.$
This proves that $P_{\sigma}$ is strongly irreducible. ∎
We can now state a result for the discrete case, where $\sigma$ is any feasible policy. The result is immediate from Lemma 3.4 and Theorem 3.3.
Corollary 3.5.
If $P_{\sigma}$ is discretely irreducible, then (E1)(E3) are equivalent.
3.4. The Significance of Irreducibility
In this section, we show that the irreducibility assumptions used in Theorems 3.3 cannot be dropped: without irreducibility, (E1)(E3) are not generally equivalent. To show this, we consider a two-state MDP with $\mathsf{X}=\{1,2\}$ and $\mathsf{A}=\{1,2\}$. The feasible coprrespondence is defined by $\Gamma(1)=\{1,2\}$ and $\Gamma(2)=\{2\}$. The reward function is defined by $r(i,j)=r_{ij}$ with
$\begin{pmatrix}r_{11}&r_{12}\\ r_{21}&r_{22}\end{pmatrix}=\begin{pmatrix}0&1\\ 0&2\end{pmatrix}.$
We set $\beta=0.9$. The transition probabilities $P(x,a,x^{\prime})$ are given by
$P(1,1,\cdot)=(1,0),\quad P(1,2,\cdot)=(1,0),\quad P(2,1,\cdot)=(0,1),\quad P(2,2,\cdot)=(0,1).$
By the definition of the feasible correspondence $\Gamma$, there are only two feasible policies $\Sigma=\{\sigma,\pi\}$, where $\sigma(x)=2$ and $\pi(x)=x$ for all $x\in\mathsf{X}$. The transition probabilities following the two policies are given by
$P_{\sigma}=P_{\pi}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
Now we compute the lifetime value functions for the optimal $\sigma$ and $\pi$. For policy $\sigma$, we have
$r_{\sigma}=(r_{12},r_{22})=(1,2),\quad v_{\sigma}=(I-\beta P_{\sigma})^{-1}r_{\sigma}=(10,20).$ (9)
For $\pi$, we have
$r_{\pi}=(r_{11},r_{22})=(0,2),\quad v_{\pi}=(I-\beta P_{\pi})^{-1}r_{\pi}=(0,20).$ (10)
Since, $v^{*}(x)\coloneq\sup_{s\in\Sigma}v_{s}(x)=v_{\sigma}(x)$, $\sigma$ is an optimal policy. On the other hand, $v_{\pi}(2)=v^{*}(2)$, but $v_{\pi}(1)<v^{*}(1)$, indicating that optimality in one state does not guarantee global optimality. This shows that irreducibility cannot be dropped from the statement of Theorem 3.3.
4. Topological Conditions
The discussion in Section 3.4 shows that strong irreducibility cannot be dropped without either (a) weakening the conclusions of Theorems 3.3, or (b) adding some side conditions. In this section, we investigate both scenarios. In particular, we show that
  1. (a)
    even when irreducibility fails, optimality can pass across some states under a continuity condition and a type of “local irreducibility,” and
  2. (b)
    when seeking the full global conclusions of Theorem 3.3, we can drop strong irreducibility if we assume a weaker form of irreducibility and pair it with continuity.
The first topic is treated in Section 4.1. The second is treated in Sections 4.2 and 4.3.
4.1. Reachable States
To begin, we return to the general MDP setting from Section 2.2, where $\mathsf{X}$ and $\mathsf{A}$ are arbitrary metric spaces. Letting $Q$ be any stochastic kernel on $\mathsf{X}$, a point $y\in\mathsf{X}$ is called $Q$-reachable from $x\in\mathsf{X}$ when, for each open neighborhood $G$ of $y$, there exists an $n\in\mathbbm{N}$ with $Q^{n}(x,G)>0$.
Theorem 4.1.
Let $\sigma$ be any continuous policy. If $v_{\sigma}(x)=v^{*}(x)$ and $y$ is $P_{\sigma}$-reachable from $x$, then $v_{\sigma}(y)=v^{*}(y)$.
Proof.
As a preliminary step, we show that $h\coloneq v^{*}-v_{\sigma}$ is continuous under the stated assumptions. Since $\sigma$ is continuous, our conditions on $(r,\Gamma,\beta,P)$ imply that the mappings $x\mapsto\int v(x^{\prime})P(x,\sigma(x),dx^{\prime})$ and $x\mapsto r(x,\sigma(x))$ are continuous on $\mathsf{X}$ whenever $v\in bc\mathsf{X}$. This implies that $T_{\sigma}$ is invariant on $bc\mathsf{X}$. Moreover, $bc\mathsf{X}$ is a closed subset of the complete metric space $bm\mathsf{X}$ under the supremum norm (since uniform limits of continuous functions are continuous). In addition, given $v,w\in bm\mathsf{X}$, we have
$\|T_{\sigma}v-T_{\sigma}w\|$ $\leqslant\beta\sup_{x\in\mathsf{X}}\int|v(x^{\prime})-w(x^{\prime})|P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime})$
$\leqslant\beta\sup_{x\in\mathsf{X}}\int\|v-w\|P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime})=\beta\|v-w\|.$
Since $\beta<1$, the contraction mapping theorem implies that $T_{\sigma}^{n}w\to v_{\sigma}$ for every $w\in bm\mathsf{X}$. If we now fix $w\in bc\mathsf{X}$ and use the fact that $T_{\sigma}$ is invariant on this set, we obtain a sequence $(T^{n}w)_{n\in\mathbbm{N}}$ converging to $v_{\sigma}$ and entirely contained in $bc\mathsf{X}$. As $bc\mathsf{X}$ is closed in $bm\mathsf{X}$, this implies that $v_{\sigma}$ is in $bc\mathsf{X}$. In particular, $v_{\sigma}$ is continuous. As $v^{*}$ is also continuous (see Proposition 2.2), we see that $h$ is continuous.
Now fix $x\in\mathsf{X}$. Seeking a contradiction, we suppose that $y$ is $P_{\sigma}$-reachable from $x$ and yet $h$ obeys $h(y)>0$. By this continuity and $h(y)>0$, there exists an open neighborhood $G$ of $y$ with $h>0$ on $G$. Because $y$ is $P_{\sigma}$-reachable from $x$, there exists an $n\in\mathbbm{N}$ with $P_{\sigma}^{n}(x,G)>0$. As a result, we have
$\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})\geqslant\int_{G}h(x^{\prime})P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})>0.$
But $v_{\sigma}(x)=v^{*}(x)$, so this inequality contradicts Lemma 2.3. The contradiction proves Theorem 4.1. ∎
Let us briefly consider how this translates to MDPs with finite state and action spaces. We give $\mathsf{X}$ and $\mathsf{A}$ the discrete topology, under which every set is open. In this setting, $y\in$ is $Q$-reachable from $x\in\mathsf{X}$ if and only if $y$ is $Q$-accessible from $x$. Moreover, every function from $\mathsf{X}$ to $\mathsf{A}$ is continuous. These observations lead to the next corollary.
Corollary 4.2.
If $v_{\sigma}(x)=v^{*}(x)$ and $y$ is $P_{\sigma}$-accessible from $x$, then $v_{\sigma}(y)=v^{*}(y)$
4.2. Open Set Irreducibility
The results in Section 4.1 discussed forms of “local” irreducibility and their implications. In this section, we analyze settings where these local conditions extend across the whole space and policies are continuous.
In general, a transition kernel $Q$ from $\mathsf{X}$ to itself is called open set irreducible if every $y\in\mathsf{X}$ is reachable from every $x\in\mathsf{X}$. For continuous policies that generate open set irreducible transitions, we have the following result.
Theorem 4.3.
If $P_{\sigma}$ is open set irreducible and $\sigma$ is continuous, then (E1)(E3) are equivalent.
Proof.
In view of Lemma 3.1, it suffices to show (E1) implies (E3). So fix $x\in\mathsf{X}$ and suppose that $v_{\sigma}(x)=v^{*}(x)$. For any $y\in\mathsf{X}$, open set irreducibility implies that $y$ is $P_{\sigma}$-reachable from $x$. Hence, by Theorem 4.1, we have $v_{\sigma}(y)=v^{*}(y)$. In particular, (E3) holds. ∎
4.3. $\pi$-Irreducibility
We treat one more form of irreducibility, due to its importance in the literature on Markov dynamics. In general, given a nontrivial measure $\pi$ on $(\mathsf{X},\mathscr{B})$, a transition kernel $Q$ on $\mathsf{X}$ is called $\pi$-irreducible if, for each $x\in X$ and every Borel set $B\subset X$ with $\pi(B)>0$, there exists an $n\in\mathbbm{N}$ such that $Q^{n}(x,B)\coloneq(Q^{n}\mathbbm{1}_{B})(x)>0$. (See, e.g., Meyn and Tweedie (2012).) Here, we will say that $Q$ is weakly irreducible if there exists a measure $\pi$ on $(\mathsf{X},\mathscr{B})$ such that
  1. (ii)
    $\pi$ assigns positive measure to all nonempty open sets, and
  2. (i)
    $Q$ is $\pi$-irreducible.
Lemma 4.4.
The following implications hold for any transition kernel $Q$ on $\mathsf{X}$.
  1. (a)
    If $Q$ is strongly irreducible, then $Q$ is weakly irreducible.
  2. (b)
    If $Q$ is weakly irreducible, then $Q$ is open set irreducible.
Proof.
Regarding (a), let $Q$ be strongly irreducible and let $\pi$ be any distribution on $\mathsf{X}$ such that $\pi(G)>0$ whenever $G\subset\mathsf{X}$ is open and nonempty.111 Such a measure exists in many settings, such as when $\mathsf{X}$ is a locally compact topological group – in which case we can take $\pi$ to be the Haar measure. In many applications, $\mathsf{X}$ will be a subset of $\mathbbm{R}^{n}$ and $\pi$ will be Lebesgue measure. Fix $B\in\mathscr{B}$ with $\pi(B)>0$ and fix $x\in B$. We recall from Lemma 2.1 that $\delta_{x}$ is a nonzero element of the dual space $b\mathsf{X}^{\prime}$. Also, $B$ is not the empty set because $\pi(B)>0$, so $\mathbbm{1}_{B}$ is a nonzero element of $b\mathsf{X}$. Hence, by strong irreducibility, there exists an $n\in\mathbbm{N}$ with $\left\langle{\delta_{x},Q^{n}\mathbbm{1}_{B}}\right\rangle>0$. We can rewrite this as $Q^{n}(x,B)=(Q^{n}\mathbbm{1}_{B})(x)>0$. This proves that $Q$ is $\pi$-irreducible. We conclude that $Q$ is weakly irreducible.
Regarding (b), let $Q$ be weakly irreducible and let $\pi$ be the measure in (i)–(ii) of the definition of weak irreducibility. Pick any $x,y\in\mathsf{X}$ and let $G$ be any open neighborhood of $y$. By (i), we have $\pi(G)>0$. By (ii), we can find an $m\in\mathbbm{N}$ with $P^{m}(x,G)>0$. Hence $y$ is reachable from $x$. Since $x$ and $y$ were chosen arbitrarily, we conclude that $Q$ is open set irreducible. ∎
Now we state a result for the weakly irreducible case. In the statement $\sigma$ is any feasible policy.
Theorem 4.5.
If $P_{\sigma}$ is weakly irreducible and $\sigma$ is continuous, then (E1)(E3) are equivalent.
Proof.
In view of Lemma 3.1, it suffices to show (E1) implies (E3). This is true by Theorem 4.3 and Lemma 4.4. ∎
5. Application: Optimal Saving with Stochastic Returns
In this section, we examine an optimal savings problem with stochastic returns on wealth. Our aim is to illustrate the theoretical results stated above. In the problem, an agent seeks to maximize lifetime utility by choosing an optimal consumption plan. The evolution of wealth is governed by the equation
$W_{t+1}=\eta_{t+1}(W_{t}-C_{t})+Y_{t+1},\qquad t=0,1,\ldots,$ (11)
where $W_{t}\in\mathbbm{R}_{+}$ is time-$t$ wealth, $C_{t}$ is the current-period consumption, $Y_{t+1}$ is the next-period labor income, and $\eta_{t+1}$ represents the stochastic return on savings. The sequences $(Y_{t})$ and $(\eta_{t})$ are IID with distributions $\varphi$ and $\psi$ respectively. For now we assume that both of these distributions have full support on $\mathbbm{R}_{+}$. To simplify the notation, we use $w^{\prime}=\eta^{\prime}(w-c)+y^{\prime}$ to denote the evolution of wealth.
We formulate this problem as an MDP. The state space is $\mathbbm{R}_{+}$ and the set of feasible actions at wealth level $w$ is $\Gamma(w)=\{c\in\mathbbm{R}_{+}:c\leqslant w\}$. A feasible policy in this setup is a Borel measurable function $\sigma$ from $\mathbbm{R}_{+}$ to itself satisfying $\sigma(w)\leqslant w$ for all $w\in\mathbbm{R}_{+}$. The reward function is $r(w,c)\coloneq u(c)$, where $u(c)$ is the utility derived from consumption and $u$ is continuous and strictly concave on $\mathbbm{R}_{+}$.
The transition kernel $P(w,c,\mathop{}\!\mathrm{d})$ is given by
$P(w,c,B)=\int\mathbbm{1}_{B}(\eta^{\prime}(w-c)+y^{\prime})\psi(d\eta^{\prime})\varphi(dy^{\prime}),$ (12)
where $0\leqslant c\leqslant w$ and $B$ is a Borel set in $\mathbbm{R}_{+}$. Given $\sigma\in\Sigma$, the corresponding policy operator $T_{\sigma}$ is given by
$(T_{\sigma}v)(w)=u(\sigma(w))+\beta(P_{\sigma}\,v)(w)\coloneq u(\sigma(w))+\beta\int v(w^{\prime})P(w,\sigma(w),\mathop{}\!\mathrm{d}w^{\prime}),$
where $\beta\in(0,1)$ is the discount factor. Using this setup, we have the following result:
Lemma 5.1.
Under the stated assumptions, the optimal savings transition kernel $P_{\sigma}$ is open set irreducible for every $\sigma\in\Sigma$.
Proof.
Let $\Sigma$ be the set of feasible policies. Fix $w\in\mathbbm{R}_{+}$ and an open set $B\subseteq\mathbbm{R}_{+}$ with $\pi(B)>0$. Let $\alpha=w-\sigma(w)$. By a change of variable, we obtain
[…]
We treat the inner integral first. Since $\eta^{\prime}$ and $y^{\prime}$ are independent random variables on $\mathbbm{R}_{+}$ with strictly positive densities $\varphi$ and $\psi$ respectively and $w^{\prime}=\alpha\eta^{\prime}+y^{\prime}$, we have the probability density function $f$ of $w^{\prime}$ at any point $w^{\prime}>0$ is given by the convolution
[…]
as in the inner integral. Since $w^{\prime}=\eta^{\prime}(w-c)+y^{\prime}$ and $y^{\prime}>0$, $w^{\prime}-\alpha\eta^{\prime}>0$, which implies that $0<\eta^{\prime}<\frac{w^{\prime}}{\alpha}$. Therefore, we have
[…]
Let
$B(w,c,v)=u(c)+\beta\int\int v(\eta^{\prime}(w-c)+y^{\prime})\psi(d\eta^{\prime})\varphi(dy^{\prime}).$
Since $u$ is strictly concave, the map $c\mapsto B(w,c,v)$ is strictly concave whenever $v$ is concave on $\mathbbm{R}_{+}$. One can also show that $v^{*}$ is concave on $\mathbbm{R}_{+}$. Combining these facts with the Bellman equation, it is straightforward to show that the optimal policy is both unique and continuous. We record this in the proposition below. More details on the arguments can be found in Chapter 12 of Stachurski (2022).
Proposition 5.2.
Under the assumptions stated above, the optimal policy of the optimal savings model is unique and continuous on $\mathbbm{R}_{+}$.
5.1. Computation
Given the open set irreducibility of the transition kernel at any feasible policy stated in Lemma 5.1, Theorem 4.3 implies that we can compute a (globally) optimal policy by maximizing $v_{\sigma}(w)$ at any fixed $w\in\mathbbm{R}_{+}$. We now explore this result in a computational experiment, where the maximization of $v_{\sigma}(w)$ is based on the Deep Deterministic Policy Gradient (DDPG) algorithm of Lillicrap et al. (2019). The DDPG algorithm uses an actor-critic (AC) method approach that iteratively improves both the policy guess and the value guess (Witten, 1977; Barto et al., 1983; Mnih et al., 2016; Wu et al., 2017; Sutton and Barto, 2018).
Our implementation uses neural networks to approximate both the optimal policy $\sigma$ and the value function $v$. Let $\theta,\lambda\in\mathbbm{R}^{d}$ be the network parameters that characterize the policy and value networks respectively. The policy network $\hat{\sigma}(\cdot;\theta)$ approximates $\sigma$, while the value network $\hat{v}_{\sigma}(\cdot;\lambda)$ approximates $v_{\sigma}$.
Let $N$ be the batch size of Monte Carlo samples. In each training episode, we draw two sequences of wealth levels: current wealth $(w_{i})_{i=1}^{N}$ and corresponding next-period wealth $(w_{i}^{\prime})_{i=1}^{N}$, where successive pairs $(w_{i},w_{i}^{\prime})$ evolve according to the wealth dynamics (11). The parameter updates in each episode of the training run as follows:
  1. (a)
    Given value parameter vector $\lambda$, the policy improvement updates the policy network parameters $\theta$ to maximize the expected return
    $J(\theta)=\frac{1}{N}\sum_{i=1}^{N}u(\hat{\sigma}(w_{i};\theta))+\beta\hat{v}_{\sigma}(w^{\prime}_{i};\lambda).$ (13)
  2. (b)
    Given policy network parameter vector $\theta$, the policy evaluation updates the value network parameters $\lambda$ to minimize the loss function
    $L(\lambda)=\frac{1}{N}\sum_{i=1}^{N}\bigg{[}\hat{v}_{\sigma}(w_{i};\lambda)-\underbrace{(u(\hat{\sigma}(w_{i};\theta))+\beta\hat{v}_{\sigma}(w^{\prime}_{i};\lambda))}_{\text{target value}}\bigg{]}^{2}.$ (14)
Full details of the algorithm are given in Algorithm 1 in Appendix. The policy network can be viewed as an actor who estimates the policy to maximize expected lifetime return, while the value network is a critic who evaluates the policy and guides the actor by providing the continuation value. The algorithm iteratively updates $\theta$ and $\lambda$ to find the optimal policy guided by a constantly improving value function.
At each iteration, the policy network aims to obtain a policy that maximizes an estimate of the $\sigma$-value function as provided by
$\hat{v}_{\sigma}(w;\lambda)=r_{\hat{\sigma}}(w)+\beta\mathbbm{E}_{w^{\prime}}[\hat{v}_{\sigma}(w^{\prime};\lambda)],$ (15)
where $r_{\hat{\sigma}}(w)=u(\hat{\sigma}(w;\theta))$. Through recursive substitution of the approximated $\sigma-$value function, we obtain
$\hat{v}_{\sigma}(w;\lambda)$ $=r_{\hat{\sigma}}(w)+\beta\mathbbm{E}_{w^{\prime}}[r_{\hat{\sigma}}(w^{\prime})+\beta\mathbbm{E}_{w^{\prime\prime}}[\hat{v}_{\sigma}(w^{\prime\prime};\lambda)]]$
$=r_{\hat{\sigma}}(w)+\beta P_{\hat{\sigma}}r_{\hat{\sigma}}(w)+\beta^{2}(P_{\hat{\sigma}})^{2}\hat{v}_{\sigma}(w;\lambda).$ (16)
Continuing this substitution yields
$\hat{v}_{\sigma}(w;\lambda)=\sum_{t=0}^{\infty}\beta^{t}(P_{\hat{\sigma}})^{t}r_{\hat{\sigma}}(w),$ (17)
which corresponds to the lifetime value of policy $\sigma$ in equation (3) evaluated at state $w\in\mathsf{W}$. Thus, our algorithm provides a practical implementation of policy optimization at a single state $w\in\mathsf{W}$.
By Lemma 5.1, the transition kernel $P_{\sigma}$ in our optimal saving problem exhibits open set irreducibility for all feasible policies $\sigma\in\Sigma$. Moreover, Proposition 5.2 ensures that the optimal policy is unique and continuous. Together with Theorem 4.3, the optimal policy can be derived from any wealth level $w\in\mathbbm{R}_{+}$ using Algorithm 1.
Let the value function obtained by following the policy network $\hat{\sigma}$ be $v_{\hat{\sigma}}$.
To verify our theory, we benchmark the result from optimizing the policy at one state against Optimistic Policy Iteration (OPI), a variant of Value Function Iteration (VFI) that is known to converge globally to the optimal policy in this model-based setting (Sargent and Stachurski, 2025). The OPI algorithm operates on a discretized state space and serves as our ground truth comparison. Specifically, we compute $v^{*}$ by applying OPI over a fine grid of wealth levels, which yields a global approximation of the optimal value function and corresponding optimal policy $\sigma^{*}$.
In contrast, our method computes $v_{\hat{\sigma}}$ by fixing a single initial wealth level $w_{0}\in\mathsf{W}$ and maximizing the value network ouput $\hat{v}_{\sigma}(w_{0})$ over the set of policies $\sigma\in\Sigma$, parameterized by the policy network vector $\theta$. We then take the resulting policy $\hat{\sigma}$ and, holding this policy fixed, calculate the entire function $v_{\hat{\sigma}}$ (or, more correctly, the value of the function at different wealth levels on a grid) by computing the lifetime value of $\hat{\sigma}$ from alternative initial conditions. Finally, we compare $v_{\hat{\sigma}}$ with the globally optimal solution $v^{*}$.
Figure 1. $v_{\hat{\sigma}}$ and $\hat{\sigma}$ with $w=50$ against the OPI solutions.
Figure 2. Expected returns (13) and value loss (14) over training episode for the irreducible optimal saving model at $w=50$.
Figure 1 shows the result of these computations when $w_{0}=50$. The function $v_{\hat{\sigma}}$, shown in blue, closely matches the globally optimal value function $v^{*}$ computed via OPI (red dotted line). Similarly, in the second panel, the approximated policy $\hat{\sigma}$ (blue line) closely matches the optimal policy $\sigma^{*}$ (red dotted line). This convergence demonstrates that both the $\hat{\sigma}$-value function $v_{\hat{\sigma}}$ and the policy $\hat{\sigma}$ successfully recover their globally optimal counterparts $v^{*}$ and $\sigma^{*}$, respectively.
In subsequent experiments, we tested the robustness of this outcome to variation in the fixed initial condition $w_{0}$. We found that, as predicted by theory, the resulting function $v_{\hat{\sigma}}$ again closely approximates $v^{*}$, and the resulting policy $\hat{\sigma}$ again closely approximates $\sigma^{*}$. See Figure 6 in Appendix B for a visualization of one experiment.
The training dynamics, illustrated in Figure 2, show consistent improvement in both performance metrics. The expected returns (13) exhibit steady increase, while the value loss (14) decreases throughout the training episodes. The training process stabilizes over episodes. Moreover, the value loss consistently remains below the levels observed in the reducible cases presented in Figures 5(a) and 5(b) of the following section.
5.2. Reducible optimal savings MDP
Now consider a modified version of the optimal saving example where both returns and labor income are bounded:
$w^{\prime}=\eta^{\prime}(w-c)+y^{\prime},\quad\eta^{\prime}\in[\underline{\eta},\overline{\eta}],\quad y^{\prime}\in[\underline{y},\overline{y}]$ (18)
with $10<\underline{\eta}<\overline{\eta}<$ 1 and $0\leqslant\underline{y}<\overline{y}<\infty$. For $w\in\mathbbm{R}_{+}$ and a Borel set $B\subseteq\mathbbm{R}_{+}$, the stochastic kernel $P_{\sigma}$ is:
$P_{\sigma}(w,B)=\int\mathbbm{1}_{B}(\eta^{\prime}(w-\sigma(w))+y^{\prime})\psi(d\eta^{\prime})\varphi(dy^{\prime})$ (19)
where $\psi$ and $\varphi$ have support on $[\underline{\eta},\overline{\eta}]$ and $[\underline{y},\overline{y}]$ respectively. In this case, we let $\psi$ and $\varphi$ be uniform distributions.
Figure 3. The law of motion of wealth given consumption $c=0$ with $\overline{\eta}=0.8$ and $\overline{y}=8$.
The next proposition shows that this optimal saving MDP is reducible.
Proposition 5.3.
For any feasible policy $\sigma\in\Sigma$, $P_{\sigma}$ is reducible.
Proof.
Fix initial wealth $w_{0}\in\mathbbm{R}_{+}$. For any $t$-step transition, let $\alpha_{t}=w_{t}-\sigma(w_{t})$ be the savings at step $t$. Then for any $\sigma\in\Sigma$
$w_{t+1}\leqslant\overline{\eta}\alpha_{t}+\overline{y}\leqslant\overline{\eta}w_{t}+\overline{y}.$
Iterating this inequality $n$ times from $w_{0}$
$w_{n}\leqslant\overline{\eta}^{n}w_{0}+\overline{y}(1+\overline{\eta}+\dots+\overline{\eta}^{n-1})=\overline{\eta}^{n}w_{0}+\overline{y}\frac{1-\overline{\eta}^{n}}{1-\overline{\eta}}.$
Hence, there exists an $M\in\mathbbm{R}_{+}$ such that
$w_{n}<\overline{\eta}w_{0}+\overline{y}\frac{1}{1-\overline{\eta}}<M\quad\forall n\in\mathbbm{N}$
Let $B=(M,\infty)$. Then $\pi(B)>0$, and $P^{n}_{\sigma}(w_{0},B)=0$ for all $n\in\mathbbm{N}$ as $w_{n}$ is bounded above by $M$ for all $n\in\mathbbm{N}$. ∎
The failure of irreducibility can be understood through the dynamics illustrated in Figure 3. When starting from any initial wealth $w\in\mathsf{W}$, the law of motion restricts the wealth process to converge towards a steady state. This convergence prevents the process from exploring the entire state space. The figure specifically demonstrates the limited range of wealth levels that can be reached from a given initial state when following the policy $\sigma(x)=0$.
Figure 4. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $w=1$ and $w=50$ against the OPI solutions with $[\underline{\eta},\overline{\eta}]=[0.5,0.8]$ and $[\underline{y},\overline{y}]=[1,8]$.
(a) $w=1$
(b) $w=50$
Figure 5. Expected returns (13) and value loss (14) over training episodes for the reducible optimal saving model.
For reducible MDPs, Algorithm 1 no longer guarantees convergence to the optimal policy when optimized at a single initial state $w\in\mathsf{W}$. This limitation is clearly demonstrated in Figure 4, which shows significant discrepancies between the algorithm’s output and the OPI solution. Figure 5 confirms this limitation, showing that extended training episodes fail to improve the policy’s performance.
Figure 4 also provides empirical support for Theorem 4.1. The policy network achieves near-optimal performance at lower wealth levels because these levels are reachable from the initial state under the learned policy $\hat{\sigma}$. However, the performance deteriorates for wealth levels that are not reachable from the initial state given the policy $\hat{\sigma}$. We note the reachable range of wealth level in the graph is smaller than in Figure 3 because $\hat{\sigma}(x)>0$ for all $x\in\mathsf{W}$.
6. Extensions and Future Work
Using MDP optimality results from Bäuerle and Rieder (2011) or Bertsekas (2022), it should be possible to extend our results to the case of unbounded rewards by replacing the ordinary supremum norm on $b\mathsf{X}$ with a weighted supremum norm. Also, while our results have focused on standard MDPs with constant discount factors, one useful variation of this model is MDPs with state-dependent discount factors, so that $\beta$ becomes a map from $\mathsf{X}$ to $\mathbbm{R}_{+}$ (see, e.g., Stachurski and Zhang (2021)). We conjecture that similar results will be available under suitable stability and irreducibility assumptions. The ideas are left for future research.
It seems likely that results similar to Theorem 3.3 will be valid for some continuous time MDPs, as well as at least some of the nonstandard discrete time dynamic programs discussed in Bertsekas (2022) and Sargent and Stachurski (2025). These topics are also left for future work.
Appendix A Algorithm
Input: Model parameters $(\beta,\gamma)$, distributions $(\psi,\varphi)$, episodes $K$, batch size $N$
Output: Policy network $\hat{\sigma}$, Value network $\hat{v}_{\sigma}$
1
2Initialize policy network $\hat{\sigma}(\cdot;\theta)$ and value network $\hat{v}_{\sigma}(\cdot;\lambda)$;
3 Initialize policy optimizer with learning rate $\alpha_{\theta}$ and value optimizer with learning rate $\alpha_{\lambda}$;
4 Initialize initial states in batch $w\in\mathbbm{R}^{N}$;
5
6for episode $k=1$ to $K$ do
       $\tilde{w}\leftarrow(w-w_{\min})/(w_{\max}-w_{\min})$ ;
        // Normalize states
7
      $c\leftarrow\hat{\sigma}(\tilde{w};\theta)$ ;
        // Current policy
       $v\leftarrow\hat{v}_{\sigma}(\tilde{w};\lambda)$ ;
        // Value estimate
       $r\leftarrow u(c)$ ;
        // Current reward
8
      $\eta\sim\psi$, $y\sim\varphi$ ;
        // Sample shocks
       $w^{\prime}\leftarrow\eta(w-c)+y$ ;
        // Next state
9       $\tilde{w}^{\prime}\leftarrow(w^{\prime}-w_{\min})/(w_{\max}-w_{\min})$ ;
10
      $v^{\prime}\leftarrow\hat{v}_{\sigma}(\tilde{w}^{\prime};\lambda)$ ;
        // Next state value
       $v_{\text{target}}\leftarrow r+\beta v^{\prime}$ ;
        // Bellman target
11
      $L\leftarrow\frac{1}{N}\sum_{i=1}^{N}(v_{i}-v_{\text{target},i})^{2}$ ;
        // Bellman error
       $\lambda\leftarrow\lambda-\alpha_{\lambda}\nabla_{\lambda}L$ ;
        // Value update
12
      $J\leftarrow\frac{1}{N}\sum_{i=1}^{N}(r_{i}+\beta\hat{v}_{\sigma}(\tilde{w}^{\prime}_{i};\lambda))$ ;
        // Expected return
       $\theta\leftarrow\theta+\alpha_{\theta}\nabla_{\theta}J$ ;
        // Policy improvement
13
      $w\leftarrow w^{\prime}$ ;
        // State evolution
14
15 end for
16return $\hat{\sigma},\hat{v}_{\sigma}$;
Algorithm 1 Deep Policy Value Iteration for Optimal Savings
Appendix B Irreducible Optimal Savings MDP at $w=1$
Figure 6. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $w=1$ against the OPI solutions.
Figure 7. Expected returns (13) and value loss (14) over training episode for the irreducible optimal saving model at $w=1$.