### v3

# Dynamic Programming: From Local Optimality to Global Optimality

Abstract.
In the theory of dynamic programming, an optimal policy is a policy whose lifetime value dominates that of all other policies from every possible initial condition in the state space. This raises a natural question: when does optimality from a single state imply optimality from every state? Working in a general setting, we provide sufficient conditions for this property that relate to reachability and irreducibility. Our results have significant implications for modern policy-based algorithms used to solve large-scale dynamic programs. We illustrate our findings by applying them to an optimal savings problem via an algorithm that implements gradient ascent in a policy space constructed from neural networks.
[…]
1. Introduction
Dynamic programming (DP) is a major branch of optimization theory with a vast range of applications. Within economics alone, applications extend from monetary and fiscal policy to asset pricing, unemployment, firm investment, wealth dynamics, commodity pricing, sovereign default, the division of labor, natural resource extraction, human capital accumulation, retirement decisions, portfolio choice, and dynamic pricing. Dynamic programs that include uncertainty are often called Markov decision processes (MDPs) and the theory of such processes has been extensively developed (see, e.g., Bäuerle and Rieder (2011), Hernández-Lerma and Lasserre (2012), Bertsekas (2012), or Bertsekas (2022)).
Much of the recent surge in interest in MDPs has been fueled by artificial intelligence and reinforcement learning (see, e.g., Bertsekas (2021) or Kochenderfer et al. (2022)). In recent years, researchers solving large-scale MDPs have moved away from value-based methods and towards policy-based methods, such as policy gradient ascent (see, e.g., Sutton et al. (1999), Lan et al. (2023), Kumar et al. (2023), Murphy (2024)). These policy-based methods seek to maximize a real-valued objective, such as
$m(\sigma)\coloneq\int v_{\sigma}(x)\rho(\mathop{}\!\mathrm{d}x)\qquad(\sigma\in\Sigma).$ (1)
Here $\sigma$ is a policy for a given MDP, mapping some state space $\mathsf{X}$ into a specified action space, $v_{\sigma}(x)$ represents the lifetime value of following the fixed policy $\sigma$, conditional on initial state $x$, and $\rho$ is a given “initial distribution.” In practice, each $\sigma$ is typically represented by a neural network. Policy-based methods often handle large problems with continuous state and action spaces more efficiently than traditional value-based methods such as value function iteration (VFI). The framework has led to numerous successful algorithms for solving complex decision-making problems, including Trust Region Policy Optimization (Schulman et al., 2015), Asynchronous Advantage Actor-Critic (Mnih et al., 2016), and Proximal Policy Optimization (Schulman et al., 2017).
Despite these successes, there is one significant disadvantage of policy gradient methods: unlike traditional DP algorithms, these methods are not guaranteed to converge to an optimal policy, even in relatively straightforward settings. To understand the issues at hand, recall that an optimal policy is a $\sigma\in\Sigma$ such that $v_{\sigma}(x)=\max_{s\in\Sigma}v_{s}(x)$ for every $x\in\mathsf{X}$; that is, a policy $\sigma$ such that following this policy in every period leads to maximum lifetime value from every initial state $x$. Even if one attains a global maximum in (1), with maximization over all $\sigma\in\Sigma$, there is no guarantee that the resulting policy will be an optimal policy. Indeed, the result depends in general on the initial distribution $\rho$, which is not one of the primitives of the DP problem.
Regarding this issue, the prevailing view is that relative optimality depends on the initial distribution $\rho$ having large support. For example, in a major study of policy gradient methods, Bhandari and Russo (2024) state that “Policy gradient methods have poor convergence properties if applied without an exploratory initial distribution” (Bhandari and Russo, 2024, p. 1910).
One disadvantage of choosing $\rho$ with large support is that computing the integral (1) becomes computationally expensive, since the function $v_{\sigma}(x)$ is expensive to evaluate. Another is that, in general, this integral cannot be computed exactly, and, moreover, the implications of a relatively poor approximation to the integral for optimality theory are not well-known. A third issue is that existing conditions for “sufficiently large supports” are difficult to test in practice, making it hard to understand whether or not the relevant theory applies. (We return to this last point below.)
These observations motivate the following question: Does there exist a large and important class of models for which it is permissible to choose $\rho$ with small support, while still maintaining the property that maximization of this criterion over all $\sigma\in\Sigma$ yields a (globally) optimal policy? At the extreme, when can $\rho$ be concentrated at a single point $x$, so that the maximization criterion $m(\sigma)$ from (1) is just $v_{\sigma}(x)$? In other words, when does optimality at a single state imply optimality at every state?
We show that, for standard MDPs on general state spaces, strong irreducibility of the Markov dynamics generated by $\sigma$ is sufficient for this property. Specifically, if a policy is optimal at a single state and has a strongly irreducible transition kernel, then this optimality propagates throughout the entire state space, making the policy globally optimal. (Here “strong irreducibility” refers to irreducibility of positive operators under the standard Riesz space definition.) In addition, we show that weaker forms of irreducibility are also sufficient for this result when $\sigma$ is also continuous.
Since gradient policy methods are commonly applied to problems with continuous state and action spaces, we avoid discreteness restrictions. In particular, for the MDPs that we consider, the state and action spaces can be arbitrary metric spaces. We also avoid placing restrictions on the class of MDPs under consideration, adopting only mild regularity conditions that imply existence of solutions. Focusing on problems where solutions exist is natural for this line of research, since we wish to examine conditions under which local optima imply global optima.
Papers examining theoretical properties of policy gradient methods have some connection to our work. For example, Bhandari and Russo (2024) examine when policy gradient ascent (actually decent) yields a globally optimal policy. To step from local to global optimality, they restrict the classes of MDPs under consideration and require that the initial distribution $\rho$ dominates the discounted state occupancy measure under the optimal policy (i.e, the measure is absolutely continuous with respect to $\rho$). This condition is difficult to verify in practice, since the optimal policy is a priori unknown (and, in fact, the object of the whole computational procedure). Here we avoid both large support restrictions on $\rho$ and explicit restrictions on the optimal policy. (At the same time, Bhandari and Russo (2024) supply many valuable new results concerning the policy gradient algorithm, and we make no direct contribution to this analysis.)
Other papers that examine theoretical properties of gradient policy methods include Khodadadian et al. (2021), Agarwal et al. (2021), and Xiao (2022). However, in these papers, the focus is on proving the convergence of $\int v_{\sigma}\mathop{}\!\mathrm{d}\rho$ to its maximal value, rather than obtaining global convergence from local convergence (as we do here). At the same time, these papers provide valuable rates of convergence for specific algorithms, which we do not discuss. A related line of research focuses on average-optimal policies in finite-state MDPs by leveraging specific state space structures under some policies. This includes exploring unichain, multichain, communicating, and weakly communicating MDPs to study algorithmic convergence (Bartlett and Tewari, 2009; Puterman, 2014). Our results focus on the continuous state setting, where gradient policy methods are usually applied.
In an application connected to the income fluctuation problem, we examine policy gradient methods under different irreducibility settings using neural network representations of policy functions. In this we follow a recent trend in economics towards solving large scale dynamic problems using neural networks to represent equilibrium objects (see, e.g., Maliar et al. (2021) or Friedl et al. (2023)). However, these papers do not examine the link between local and global optimality. In fact they do not use the class of policy gradient methods described above, focusing instead on minimizing Euler or Bellman residuals.
The paper is structured as follows. Section 2 provides background on MDPs. Section 3 states our main results in a general setting. Section 4 examines how the irreducibility condition from Section 3 can be weakened while still obtaining some transmission of optimality across states. Section 5.1 illustrates our theoretical results in the context of a benchmark optimal savings problem. Section 6 outlines avenues for future work.
2. Markov Decision Processes
In this section, we review essential properties of Markov decision processes and state a technical lemma that will be applied in our main results.
2.1. Preliminaries
[…]
A transition kernel on $\mathsf{X}$ is a function $Q$ from $\mathsf{X}\times\mathscr{B}$ to $[0,1]$ such that $x\mapsto Q(x,B)$ is Borel measurable for all $B\in\mathscr{B}$ and $B\mapsto Q(x,B)$ is a Borel probability measure for all $x\in\mathsf{X}$. To any such kernel $Q$ we associate a bounded linear operator on $b\mathsf{X}$, often referred to as its Markov operator and also denoted by $Q$, via
$f\mapsto Qf,\qquad(Qf)(x)=\int f(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime}).$ (2)
Typically, $(Qf)(x)$ represents the expectation of $f(X_{t+1})$ given that $X_{t}=x$ and $X_{t+1}$ is drawn from $Q(x,\mathop{}\!\mathrm{d}x^{\prime})$.
As usual, the positive cone of $b\mathsf{X}$, denoted here by $b\mathsf{X}_{+}$, is all $v\in b\mathsf{X}$ with $v\geqslant 0$. Let $b\mathsf{X}^{\prime}$ be the dual space of $b\mathsf{X}$ and let $b\mathsf{X}^{\prime}_{+}$ to be the positive cone of $b\mathsf{X}^{\prime}$. The set $b\mathsf{X}^{\prime}_{+}$ contains, among other objects, the set $\mathcal{D}(\mathsf{X})$ of Borel probability measures on $\mathsf{X}$. For simplicity, elements of $\mathcal{D}(\mathsf{X})$ are referred to as distributions. For $\rho\in\mathcal{D}(\mathsf{X})$ and $f\in b\mathsf{X}$ we set
$\left\langle{f,\rho}\right\rangle\coloneq\int f\mathop{}\!\mathrm{d}\rho.$
For each $x\in\mathsf{X}$, the point evaluation functional generated by $x$ is the map $\delta_{x}$ that sends $w\in b\mathsf{X}$ into $w(x)\in\mathbbm{R}$. Below it will be convenient for us to write this in dual notation, so that $\left\langle{w,\delta_{x}}\right\rangle=w(x)$ for every $w\in b\mathsf{X}$. We will make use of the following lemma.
Lemma 2.1.
Every point evaluation functional on $b\mathsf{X}$ is a nonzero element of $b\mathsf{X}_{+}^{\prime}$.
Proof.
Fix $x\in\mathsf{X}$. Linearity of $\delta_{x}$ is obvious: given $a,b\in\mathbbm{R}$ and $v,w\in b\mathsf{X}$, we have
$\left\langle{av+bw,\delta_{x}}\right\rangle=(av+bw)(x)=av(x)+bw(x)=a\left\langle{v,\delta_{x}}\right\rangle+b\left\langle{w,\delta_{x}}\right\rangle.$
Regarding continuity, if $w_{n}\to w$ in $b\mathsf{X}$, then $w_{n}\to w$ pointwise on $\mathsf{X}$, so $\left\langle{w_{n},\delta_{x}}\right\rangle=w_{n}(x)\to w(x)=\left\langle{w,\delta_{x}}\right\rangle$. Regarding positivity, it suffices to show that $\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$ whenever $w\geqslant 0$. This clearly holds, since $w\geqslant 0$ implies $w(x)=\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$. Finally, $\delta_{x}$ is not the zero element of $b\mathsf{X}^{\prime}$ because $w=\mathbf{1}$ is in $b\mathsf{X}$ and $\left\langle{w,\delta_{x}}\right\rangle=w(x)=1\not=0$. ∎
2.2. Markov Decision Process
Let $\mathsf{X}$ and $\mathsf{A}$ be metric spaces, as in Section 2.1. A Markov decision process (MDP) with state space $\mathsf{X}$ and action space $\mathsf{A}$ is a tuple $(r,\Gamma,\beta,P)$, where $r$ is a reward function, $x\mapsto\Gamma(x)\subset\mathsf{A}$ is a feasible correspondence, $\beta\in\mathbbm{R}$ is a discount factor and $P(x,a,\mathop{}\!\mathrm{d}x^{\prime})$ is a distribution over next period states given current state $x$ and action $a$. Let $\mathsf{G}\coloneq\{(x,a)\in\mathsf{X}\times\mathsf{A}:a\in\Gamma(x)\}$. We consider a relatively standard environment, as considered in, say, Bäuerle and Rieder (2011) and Hernández-Lerma and Lasserre (2012), where
  1. (a)
    $\beta\in(0,1)$,
  2. (b)
    $\Gamma$ is nonempty, continuous, and compact-valued on $\mathsf{X}$,
  3. (c)
    $r$ is bounded and continuous on $\mathsf{G}$, and
  4. (d)
    the map $(x,a)\mapsto\int v(x^{\prime})P(x,a,dx^{\prime})$ is continuous on $\mathsf{G}$ whenever $v\in cb\mathsf{X}$.
(The case of unbounded $r$ is discussed in Section 6.)
Let $\Sigma$ denote the set of feasible policies, by which we mean all Borel measurable functions $\sigma$ mapping $\mathsf{X}$ to $\mathsf{A}$ with $\sigma(x)\in\Gamma(x)$ for all $x\in\mathsf{X}$. For each $\sigma\in\Sigma$ and $x\in\mathsf{X}$, we set
$r_{\sigma}(x)\coloneq r(x,\sigma(x))\quad\text{and}\quad P_{\sigma}(x,\mathop{}\!\mathrm{d}x^{\prime})\coloneq P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime}).$
Thus, $r_{\sigma}(x)$ is rewards at $x$ under policy $\sigma$ and $P_{\sigma}$ is the transition kernel on $\mathsf{X}$ generated by $\sigma$. Using the corresponding Markov operator $P_{\sigma}$, as defined in (2), the lifetime value of a policy $\sigma$, denoted by $v_{\sigma}$, can be expressed as
$v_{\sigma}=\sum_{t=0}^{\infty}(\beta P_{\sigma})^{t}r_{\sigma}$ (3)
(see, e.g., Puterman (2014), Theorem 6.1.1). We will use the fact that $v_{\sigma}$ is also the unique fixed point in $b\mathsf{X}$ of the policy operator $T_{\sigma}$ defined by $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$. This operator can be written more explicitly as
$(T_{\sigma}\,v)(x)=r(x,\sigma(x))+\beta\int v(x^{\prime})P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime})\qquad(v\in b\mathsf{X},\;x\in\mathsf{X}).$
(The lifetime value $v_{\sigma}$ is the unique fixed point of $T_{\sigma}$ because the spectral radius of the linear operator $\beta P_{\sigma}$ is $\beta$, so, by the geometric series lemma, $\beta<1$ implies that $v=r_{\sigma}+\beta P_{\sigma}\,v$ has the unique solution given by the right-hand side of (3).) Iterating on the definition $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$, we find that
$T_{\sigma}^{n}v=r_{\sigma}+\beta P_{\sigma}r_{\sigma}+\cdots+(\beta P_{\sigma})^{n-1}r_{\sigma}+(\beta P_{\sigma})^{n}v\quad\text{for all $n\in\mathbbm{N}$}.$ (4)
This expression will be useful in the theory below.
The value function is denoted $v^{*}$ and defined at each $x\in\mathsf{X}$ by $v^{*}(x)\coloneq\sup_{\sigma\in\Sigma}v_{\sigma}(x)$. A policy $\sigma$ is called optimal if $v_{\sigma}(x)=v^{*}(x)$ for all $x\in\mathsf{X}$.
We define the Bellman operator by
$(Tv)(x)=\max_{a\in\Gamma(x)}\left\{r(x,a)+\beta\int v(x^{\prime})P(x,a,\mathop{}\!\mathrm{d}x^{\prime})\right\}\qquad(v\in b\mathsf{X},\;x\in\mathsf{X}).$ (5)
We will use the following facts:
Proposition 2.2.
Under the stated assumptions,
  1. (a)
    the value function $v^{*}$ is the unique fixed point of the Bellman operator in $b\mathsf{X}$,
  2. (b)
    the value function $v^{*}$ is well-defined and contained in $cb\mathsf{X}$, and
  3. (c)
    at least one optimal policy exists.
Proof.
See Bäuerle and Rieder (2011) or Hernández-Lerma and Lasserre (2012). ∎
We also make use of the following technical lemma, which shows one implication of local optimality at some given $x\in\mathsf{X}$.
Lemma 2.3.
If $\sigma\in\Sigma$ and $v_{\sigma}(x)=v^{*}(x)$, then
$\sum_{n\in\mathbbm{N}}\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})=0.$ (6)
Proof.
Fix $n\in\mathbbm{N}$. Applying the expression for $T_{\sigma}^{n}v$ from (4) twice, first with $v=v_{\sigma}$ and then with $v=v^{*}$, we get
$T_{\sigma}^{n}\,v_{\sigma}-T_{\sigma}^{n}\,v^{*}=\beta^{n}(P_{\sigma}^{n}\,v_{\sigma}-P_{\sigma}^{n}\,v^{*}).$ (7)
In addition, we have
$v_{\sigma}=T_{\sigma}^{n}\,v_{\sigma}\leqslant T_{\sigma}^{n}\,v^{*}\leqslant T^{n}\,v^{*}=v^{*}.$ (8)
In (8), the first inequality is due to the fact $T_{\sigma}$ is order preserving and $v_{\sigma}\leqslant v^{*}$, while the second follows from the fact that $T_{\sigma}\,v\leqslant Tv$ for all $v\in b\mathsf{X}$. (Since $T_{\sigma}\,v\leqslant T\,v$ for all $v$, $T_{\sigma}^{2}\,v^{*}\leqslant T_{\sigma}T\,v^{*}\leqslant T^{2}\,v^{*}$ and so $T_{\sigma}^{2}\,v^{*}\leqslant T^{2}\,v^{*}.$ By induction, the second inequality holds.) The claim in Lemma 2.3 follows from (7) and (8). To see this, fix $x\in\mathsf{X}$ with $v_{\sigma}(x)=v^{*}(x)$. From this equality and (8) we get $(T_{\sigma}^{n}\,v_{\sigma})(x)=(T_{\sigma}^{n}\,v^{*})(x)$. Since $\beta>0$, combining this result with (7) yields $(P_{\sigma}^{n}\,v_{\sigma})(x)=(P_{\sigma}^{n}\,v^{*})(x)$. Hence (6) holds. ∎
3. From Local to Global Optimality
The standard definition of optimality, which was given in Section 2.2, is global in nature, since it concerns the lifetime value of the policy at every $x\in\mathsf{X}$. We seek conditions under which local optimality implies global optimality. In particular, we seek conditions under which the following three statements are equivalent:
  1. (E1)
    there exists an $x\in\mathsf{X}$ such that $v_{s}(x)\leqslant v_{\sigma}(x)$ for all $s\in\Sigma$,
  2. (E2)
    there exists a $\rho\in\mathcal{D}(\mathsf{X})$ such that $\left\langle{v_{s},\rho}\right\rangle\leqslant\left\langle{v_{\sigma},\rho}\right\rangle$ for all $s\in\Sigma$,
  3. (E3)
    $\sigma$ is an optimal policy.
3.1. Preliminary Results
We first note that, in the MDP set up we have described, (E1) and (E2) are always equivalent, as the next lemma shows.
Lemma 3.1.
For all $\sigma\in\Sigma$, the statements (E1) and (E2) are equivalent.
Proof.
To show (E1) implies (E2), assume (E1) and fix $x\in\mathsf{X}$ with $v_{\sigma}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$. Let $\delta_{x}$ be the point evaluation functional generated by $x$. Since $\delta_{x}\in\mathcal{D}(\mathsf{X})$ and $\left\langle{v_{\sigma},\delta_{x}}\right\rangle=v_{\sigma}(x)\geqslant v_{s}(x)=\left\langle{v_{s},\delta_{x}}\right\rangle$ for all $s\in\Sigma$, so (E2) holds. To show (E2) implies (E1), fix $\rho\in\mathcal{D}(\mathsf{X})$ with $\left\langle{v_{\sigma},\rho}\right\rangle\geqslant\left\langle{v_{s},\rho}\right\rangle$ for all $s\in\Sigma$. Since at least one $s\in\Sigma$ is optimal (Proposition 2.2), this yields $\left\langle{v_{\sigma},\rho}\right\rangle\geqslant\left\langle{v^{*},\rho}\right\rangle$. Now suppose to the contrary that for each $x\in\mathsf{X},$ we can find a $\tau\in\Sigma$ such that $v_{\tau}(x)>v_{\sigma}(x)$. Since $v^{*}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$ and $x\in\mathsf{X}$, we have $v^{*}(x)>v_{\sigma}(x)$ for all $x\in\mathsf{X}$. Hence $\left\langle{v^{*},\rho}\right\rangle>\left\langle{v_{\sigma},\rho}\right\rangle$. This contradiction proves (E1). ∎
Our second preliminary observation suggests one way to obtain (E3) from (E2).
Lemma 3.2.
If $\left\langle{v_{\sigma},\rho}\right\rangle=\left\langle{v^{*},\rho}\right\rangle$, then $v_{\sigma}=v^{*}$ holds $\rho$-almost everywhere.
Proof.
By definition, $v^{*}\geqslant v_{\sigma}$ on $\mathsf{X}$. If, in addition, $v^{*}>v_{\sigma}$ on a set $E$ of positive $\rho$-measure, then
$\int(v^{*}-v_{\sigma})\mathop{}\!\mathrm{d}\rho\geqslant\int_{E}(v^{*}-v_{\sigma})\mathop{}\!\mathrm{d}\rho>0.$
This contradicts $\left\langle{v_{\sigma},\rho}\right\rangle=\left\langle{v^{*},\rho}\right\rangle$. Hence $v_{\sigma}=v^{*}$ holds $\rho$-almost everywhere. ∎
Loosely speaking, this lemma shows that maximizing the scalar performance measure $\left\langle{v_{\sigma},\rho}\right\rangle$ with a highly exploratory initial distribution $\rho$ promotes global optimality.
The disadvantage of this approach is that evaluating the scalar measure from a such an initial distribution is costly. For this reason, our main interest is in providing conditions under which (E1) implies (E3). Our conditions relate to irreducibility. The first condition, discussed in Section 3.2, uses a notion of irreducibility from the literature on Banach lattices (see, e.g., Zaanen (2012)), while the second, discussed in Section 4.3, uses an irreducibility concept from the Markov process literature.
3.2. Strong Irreducibility
Recall that a linear subspace $I$ of $b\mathsf{X}$ is called an ideal in $b\mathsf{X}$ when $f\in I$ and $|g|\leqslant|f|$ implies $g\in I$. An ideal $I$ is said to be invariant for a linear operator $K$ if $KI\subset I$. A linear operator $K$ from $b\mathsf{X}$ to itself is called positive when $Kf\geqslant 0$ for all $f\geqslant 0$. A positive linear operator $K$ is called irreducible if the only invariant ideals under $K$ are the trivial subspace $\{0\}$ and the whole space $b\mathsf{X}$. By Proposition 8.3 (c) of Schaefer (1974), such a $K$ is irreducible if and only if, for each nonzero $f\in b\mathsf{X}_{+}$ and each nonzero $\mu\in b\mathsf{X}^{\prime}_{+}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mu,K^{m}f}\right\rangle>0$.
In what follows, we call a transition kernel $P$ on $\mathsf{X}$ strongly irreducible if its Markov operator (see (2)) is irreducible on $b\mathsf{X}$ in the sense just defined. Here is our main result for the strongly irreducible case. In the statement, $\sigma$ is any feasible policy.
Theorem 3.3.
If $P_{\sigma}$ is strongly irreducible, then (E1)(E3) are equivalent.
For example, Theorem 3.3 tells us that, under the stated conditions, we can obtain an optimal policy by fixing an arbitrary initial state $x\in\mathsf{X}$ and maximizing the real-valued function $s\mapsto v_{s}(x)$ over $\Sigma$. Alternatively, we can fix any distribution $\rho$ and maximize $s\mapsto\left\langle{v_{s},\rho}\right\rangle$.
Proof of Theorem 3.3.
In view of Lemma 3.1, it suffices to show that (E1) and (E3) are equivalent. That (E3) implies (E1) is immediate from the definition of optimal policies. Hence we need only show that (E1) implies (E3). In line with the conditions of Theorem 3.3, we assume that $\sigma$ is a feasible policy and $P_{\sigma}$ is strongly irreducible. Using (E1), we take $x\in\mathsf{X}$ with $v_{\sigma}(x)=v^{*}(x)$.
Let $h\coloneq v^{*}-v_{\sigma}$. By the definition of $v^{*}$ we have $0\leqslant h$. We claim in addition that $h=0$. To see this, suppose to the contrary that $h$ is nonzero. In this case, by strong irreducibility, for each nonzero $\mu$ in the positive cone of $b\mathsf{X}^{\prime}$ we can find an $m\in\mathbbm{N}$ such that $\left\langle{\mu,P_{\sigma}^{m}\,h}\right\rangle>0$. Because $\delta_{x}$ is a nonzero element of the positive cone of $b\mathsf{X}^{\prime}$ (Lemma 2.1), we can set $\mu=\delta_{x}$ to obtain an $m\in\mathbbm{N}$ with $(P_{\sigma}^{m}\,h)(x)>0$. This contradicts (6), so $h=0$ holds. In other words, $v_{\sigma}=v^{*}$. This proves (E3). ∎
4. Topological Conditions
The discussion in Section 5.4 shows that strong irreducibility cannot be dropped without either (a) weakening the conclusions of Theorems 3.3, or (b) adding some side conditions. In this section, we investigate both scenarios. In particular, we show that
  1. (a)
    even when irreducibility fails, optimality can pass across some states under a continuity condition and a type of “local irreducibility,” and
  2. (b)
    when seeking the full global conclusions of Theorem 3.3, we can drop strong irreducibility if we assume a weaker form of irreducibility and pair it with continuity.
The first topic is treated in Section 4.1. The second is treated in Sections 4.2 and 4.3.
4.1. Reachable States
To begin, we return to the general MDP setting from Section 2.2, where $\mathsf{X}$ and $\mathsf{A}$ are arbitrary metric spaces. Letting $Q$ be any stochastic kernel on $\mathsf{X}$, a point $y\in\mathsf{X}$ is called $Q$-reachable from $x\in\mathsf{X}$ when, for each open neighborhood $G$ of $y$, there exists an $n\in\mathbbm{N}$ with $Q^{n}(x,G)>0$.
Lemma 4.1.
Let $\sigma$ be any continuous policy. If $v_{\sigma}(x)=v^{*}(x)$ and $y$ is $P_{\sigma}$-reachable from $x$, then $v_{\sigma}(y)=v^{*}(y)$.
Proof.
As a preliminary step, we show that $h\coloneq v^{*}-v_{\sigma}$ is continuous under the stated assumptions. Since $\sigma$ is continuous, our conditions on $(r,\Gamma,\beta,P)$ imply that the mappings $x\mapsto\int v(x^{\prime})P(x,\sigma(x),dx^{\prime})$ and $x\mapsto r(x,\sigma(x))$ are continuous on $\mathsf{X}$ whenever $v\in cb\mathsf{X}$. This implies that $T_{\sigma}$ is invariant on $cb\mathsf{X}$. Moreover, $cb\mathsf{X}$ is a closed subset of the complete metric space $b\mathsf{X}$ under the supremum norm (since uniform limits of continuous functions are continuous). In addition, given $v,w\in b\mathsf{X}$, we have
$\|T_{\sigma}v-T_{\sigma}w\|$ $\leqslant\beta\sup_{x\in\mathsf{X}}\int|v(x^{\prime})-w(x^{\prime})|P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime})$
$\leqslant\beta\sup_{x\in\mathsf{X}}\int\|v-w\|P(x,\sigma(x),\mathop{}\!\mathrm{d}x^{\prime})=\beta\|v-w\|.$
Since $\beta<1$, the contraction mapping theorem implies that $T_{\sigma}^{n}w\to v_{\sigma}$ for every $w\in b\mathsf{X}$. If we now fix $w\in cb\mathsf{X}$ and use the fact that $T_{\sigma}$ is invariant on this set, we obtain a sequence $(T^{n}w)_{n\in\mathbbm{N}}$ converging to $v_{\sigma}$ and entirely contained in $cb\mathsf{X}$. As $cb\mathsf{X}$ is closed in $b\mathsf{X}$, this implies that $v_{\sigma}$ is in $cb\mathsf{X}$. In particular, $v_{\sigma}$ is continuous. As $v^{*}$ is also continuous (see Proposition 2.2), we see that $h$ is continuous.
Now fix $x\in\mathsf{X}$. Seeking a contradiction, we suppose that $y$ is $P_{\sigma}$-reachable from $x$ and yet $h$ obeys $h(y)>0$. By this continuity and $h(y)>0$, there exists an open neighborhood $G$ of $y$ with $h>0$ on $G$. Because $y$ is $P_{\sigma}$-reachable from $x$, there exists an $n\in\mathbbm{N}$ with $P_{\sigma}^{n}(x,G)>0$. As a result, we have
$\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})\geqslant\int_{G}h(x^{\prime})P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})>0.$
But $v_{\sigma}(x)=v^{*}(x)$, so this inequality contradicts Lemma 2.3. The contradiction proves Lemma 4.1. ∎
4.2. Open Set Irreducibility
The results in Section 4.1 discussed forms of “local” irreducibility and their implications. In this section, we analyze settings where these local conditions extend across the whole space and policies are continuous.
In general, a transition kernel $Q$ from $\mathsf{X}$ to itself is called open set irreducible if every $y\in\mathsf{X}$ is reachable from every $x\in\mathsf{X}$. For continuous policies that generate open set irreducible transitions, we have the following result.
Theorem 4.2.
If $P_{\sigma}$ is open set irreducible and $\sigma$ is continuous, then (E1)(E3) are equivalent.
Proof.
In view of Lemma 3.1, it suffices to show (E1) implies (E3). So fix $x\in\mathsf{X}$ and suppose that $v_{\sigma}(x)=v^{*}(x)$. For any $y\in\mathsf{X}$, open set irreducibility implies that $y$ is $P_{\sigma}$-reachable from $x$. Hence, by Lemma 4.1, we have $v_{\sigma}(y)=v^{*}(y)$. In particular, (E3) holds. ∎
4.3. $\pi$-Irreducibility
We treat one more form of irreducibility, due to its importance in the literature on Markov dynamics. In general, given a nontrivial measure $\pi$ on $(\mathsf{X},\mathscr{B})$, a transition kernel $Q$ on $\mathsf{X}$ is called $\pi$-irreducible if, for each $x\in X$ and every Borel set $B\subset X$ with $\pi(B)>0$, there exists an $n\in\mathbbm{N}$ such that $Q^{n}(x,B)\coloneq(Q^{n}\mathbbm{1}_{B})(x)>0$. (See, e.g., Meyn and Tweedie (2012).) Here, we will say that $Q$ is weakly irreducible if there exists a measure $\pi$ on $(\mathsf{X},\mathscr{B})$ such that
  1. (ii)
    $\pi$ assigns positive measure to all nonempty open sets, and
  2. (i)
    $Q$ is $\pi$-irreducible.
Lemma 4.3.
The following implications hold for any transition kernel $Q$ on $\mathsf{X}$.
  1. (a)
    If $Q$ is strongly irreducible, then $Q$ is weakly irreducible.
  2. (b)
    If $Q$ is weakly irreducible, then $Q$ is open set irreducible.
Proof.
Regarding (a), let $Q$ be strongly irreducible and let $\pi$ be any distribution on $\mathsf{X}$ such that $\pi(G)>0$ whenever $G\subset\mathsf{X}$ is open and nonempty.111 Such a measure exists in many settings, such as when $\mathsf{X}$ is a locally compact topological group – in which case we can take $\pi$ to be the Haar measure. In many applications, $\mathsf{X}$ will be a subset of $\mathbbm{R}^{n}$ and $\pi$ will be Lebesgue measure. Fix $B\in\mathscr{B}$ with $\pi(B)>0$ and fix $x\in B$. We recall from Lemma 2.1 that $\delta_{x}$ is a nonzero element of the dual space $b\mathsf{X}^{\prime}$. Also, $B$ is not the empty set because $\pi(B)>0$, so $\mathbbm{1}_{B}$ is a nonzero element of $b\mathsf{X}$. Hence, by strong irreducibility, there exists an $n\in\mathbbm{N}$ with $\left\langle{\delta_{x},Q^{n}\mathbbm{1}_{B}}\right\rangle>0$. We can rewrite this as $Q^{n}(x,B)=(Q^{n}\mathbbm{1}_{B})(x)>0$. This proves that $Q$ is $\pi$-irreducible. We conclude that $Q$ is weakly irreducible.
Regarding (b), let $Q$ be weakly irreducible and let $\pi$ be the measure in (i)–(ii) of the definition of weak irreducibility. Pick any $x,y\in\mathsf{X}$ and let $G$ be any open neighborhood of $y$. By (i), we have $\pi(G)>0$. By (ii), we can find an $m\in\mathbbm{N}$ with $P^{m}(x,G)>0$. Hence $y$ is reachable from $x$. Since $x$ and $y$ were chosen arbitrarily, we conclude that $Q$ is open set irreducible. ∎
Now we state a result for the weakly irreducible case. In the statement, $\sigma$ is any feasible policy.
Theorem 4.4.
If $P_{\sigma}$ is weakly irreducible and $\sigma$ is continuous, then (E1)(E3) are equivalent.
Proof.
In view of Lemma 3.1, it suffices to show (E1) implies (E3). This is true by Theorem 4.2 and Lemma 4.3. ∎
5. Applications
In this section, we consider three applications. The first involves an optimal savings problem with stochastic returns on wealth, and our aim is to illustrate the significance of our main theoretical results. The second application provides an extension showing how results in the paper can be applied outside the MDP framework. The third application, found in Section 5.4, demonstrates that the irreducibility assumptions in Theorems 3.3, 4.2, and 4.4 cannot be removed without changing their conclusions.
5.1. Optimal Savings
Consider an agent who seeks to maximize lifetime utility by choosing optimal savings and consumption. The evolution of wealth is governed by the equation
$w_{t+1}=\eta_{t+1}(w_{t}-c_{t})+y_{t+1},\qquad t=0,1,\ldots,$ (9)
where $w_{t}\in\mathbbm{R}_{+}$ is time-$t$ wealth, $c_{t}$ is the current-period consumption, $y_{t+1}$ is the next-period labor income, and $\eta_{t+1}$ represents the stochastic return on savings. The sequences $(y_{t})$ and $(\eta_{t})$ are IID with distributions $\varphi$ and $\psi$ respectively. For now we assume that both of these distributions have full support on $\mathbbm{R}_{+}$. To simplify the notation, we use $w^{\prime}=\eta^{\prime}(w-c)+y^{\prime}$ to denote the evolution of wealth.
We formulate this problem as an MDP. The state space is $\mathbbm{R}_{+}$ and the set of feasible actions at wealth level $w$ is $\Gamma(w)=\{c\in\mathbbm{R}_{+}:c\leqslant w\}$. A feasible policy in this setup is a Borel measurable function $\sigma$ from $\mathbbm{R}_{+}$ to itself satisfying $\sigma(w)\leqslant w$ for all $w\in\mathbbm{R}_{+}$. The reward function is $r(w,c)\coloneq u(c)$, where $u(c)$ is the utility derived from consumption and $u$ is continuous, differentiable, and strictly concave on $\mathbbm{R}_{+}$.
The transition kernel $P(w,c,\mathop{}\!\mathrm{d})$ is given by
$P(w,c,B)=\int\mathbbm{1}_{B}(\eta^{\prime}(w-c)+y^{\prime})\psi(d\eta^{\prime})\varphi(dy^{\prime}),$ (10)
where $0\leqslant c\leqslant w$ and $B$ is a Borel set in $\mathbbm{R}_{+}$. Given $\sigma\in\Sigma$, the corresponding policy operator $T_{\sigma}$ is given by
$(T_{\sigma}v)(w)=u(\sigma(w))+\beta(P_{\sigma}\,v)(w)\coloneq u(\sigma(w))+\beta\int v(w^{\prime})P(w,\sigma(w),\mathop{}\!\mathrm{d}w^{\prime}),$
where $\beta\in(0,1)$ is the discount factor. Using this setup, we have the following result:
Lemma 5.1.
Under the stated assumptions, the optimal savings transition kernel $P_{\sigma}$ is open set irreducible for every $\sigma\in\Sigma$.
Proof.
Let $\Sigma$ be the set of feasible policies. Fix $\bar{w}\in\mathbbm{R}_{+}$ and an open set $B\subseteq\mathbbm{R}_{+}$ with $\pi(B)>0$. Let $\alpha=w-\sigma(w)$. By a change of variable, we obtain
[…]
We treat the inner integral first. Since $\eta^{\prime}$ and $y^{\prime}$ are independent random variables on $\mathbbm{R}_{+}$ with strictly positive densities $\varphi$ and $\psi$ respectively and $w^{\prime}=\alpha\eta^{\prime}+y^{\prime}$, we have the probability density function $f$ of $w^{\prime}$ at any point $w^{\prime}>0$ is given by the convolution
[…]
as in the inner integral. Since $w^{\prime}=\eta^{\prime}(w-c)+y^{\prime}$ and $y^{\prime}>0$, $w^{\prime}-\alpha\eta^{\prime}>0$, which implies that $0<\eta^{\prime}<\frac{w^{\prime}}{\alpha}$. Therefore, we have
[…]
Let
$B(w,c,v)=u(c)+\beta\int\int v(\eta^{\prime}(w-c)+y^{\prime})\psi(d\eta^{\prime})\varphi(dy^{\prime}).$
Since $u$ is strictly concave, the map $c\mapsto B(w,c,v)$ is strictly concave whenever $v$ is concave on $\mathbbm{R}_{+}$. One can also show that $v^{*}$ is concave on $\mathbbm{R}_{+}$. Combining these facts with the Bellman equation, it is straightforward to show that the optimal policy is both unique and continuous. We record this in the proposition below. More details on the arguments can be found in Chapter 12 of Stachurski (2022).
Proposition 5.2.
Under the assumptions stated above, the optimal policy of the optimal savings model is unique and continuous on $\mathbbm{R}_{+}$.
Given the open set irreducibility of the transition kernel at any feasible policy stated in Lemma 5.1, Theorem 4.2 implies that we can compute a (globally) optimal policy by maximizing $v_{\sigma}(\bar{w})$ at any fixed $\bar{w}\in\mathbbm{R}_{+}$. We now explore this result in a computational experiment, where the maximization of $v_{\sigma}(\bar{w})$ is based on a simple gradient ascent algorithm with a neural network to approximate the policy function $\sigma$.
Let $\theta\in\mathbbm{R}^{d}$ be the network parameters that characterize the policy network $\hat{\sigma}(\cdot;\theta)$ with a fixed paramter size $d\in\mathbbm{N}$. We want to find the policy parameterization that maximizes the expected discounted sum of rewards defined by $u(\hat{\sigma}(w;\theta))$ starting from the initial wealth level $\bar{w}$:
$\max_{\hat{\sigma}\in\Sigma}v_{\hat{\sigma}}(\bar{w})\coloneq\max_{\hat{\sigma}\in\Sigma}\sum_{t=0}^{\infty}\beta^{t}\int u(\hat{\sigma}(w^{\prime};\theta))P^{t}_{\hat{\sigma}}(\bar{w},\mathop{}\!\mathrm{d}w^{\prime}).$
We can approximate this by Monte Carlo with large batch size $N$ and rollout steps $T$:
$\hat{v}_{\hat{\sigma}}(\bar{w})\coloneq\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\beta^{t}u(\hat{\sigma}(w_{i,t};\theta))\quad\text{with}\quad w_{i,0}=\bar{w}\text{ for }i=1,\ldots,N,$ (11)
Here $w_{i,t}$ is the wealth at time $t$ for the $i$-th sample, conditional on starting at $\bar{w}$ and following policy $\hat{\sigma}$. (In particular, we generate the rollout trajectories $(w_{i,t})_{t=0}^{T}$ for all $i=1,\ldots,N$ according to the wealth dynamics (9) with consumption at time $t$ following $\hat{\sigma}(w_{i,t};\theta)$.) The fact that $w_{i,0}=\bar{w}$ for all $i=1,\ldots,N$ means that we are optimizing lifetime value over policies while freezing the initial state $w_{0}=\bar{w}$.
During each episode, the model parameters are updated to reduce the value of the loss function
$L(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\beta^{t}u(\hat{\sigma}(w_{i,t};\theta))$ (12)
using gradient descent. It is clear that minimizing the loss is equivalent to maximizing the finite time approximation of the expected lifetime value (11).
The complete algorithm is detailed in Algorithm 1 in the Appendix. Our approach is similar to the REINFORCE algorithm (Williams, 1992), with a key distinction: since the utility function in our setup is differentiable, we can compute well-defined gradients without requiring a stochastic policy.
Several techniques could enhance the algorithm’s stability and convergence, including reinforcement learning baselines (Williams, 1992) and more advanced approaches applying actor-critic methods (Silver et al., 2014; Lillicrap et al., 2019). We deliberately avoid these extensions in this paper to maintain algorithmic clarity and simplicity, allowing us to establish a more direct connection to our theoretical results.
Figure 1. $v_{\hat{\sigma}}$ and $\hat{\sigma}$ with $\bar{w}=1$ vs OPI solutions
Figure 2. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the irreducible optimal saving model at $\bar{w}=1$
To verify our theory, we benchmark the result from optimizing the policy at one point against Optimistic Policy Iteration (OPI), a variant of Value Function Iteration (VFI) that is known to converge globally to the optimal policy in this model-based setting (Sargent and Stachurski, 2025a). The OPI algorithm operates on a discretized state space and serves as our ground truth comparison. Specifically, we compute $v^{*}$ by applying OPI over a fine grid of wealth levels, which yields a global approximation of the optimal value function and corresponding optimal policy $\sigma^{*}$.
Regarding $v_{\hat{\sigma}}$, we compute this function in two steps. First, we freeze a single initial wealth level $\bar{w}$ and minimize the loss according to (12), as described above. Second, we take the resulting policy $\hat{\sigma}$ and, holding this policy fixed, calculate the function $v_{\hat{\sigma}}$ by computing the lifetime value of $\hat{\sigma}$ from alternative initial conditions. Finally, we compare $v_{\hat{\sigma}}$ with the globally optimal solution $v^{*}$. In the experiments, we use the CRRA utility function $u(c)=c^{1-\gamma}/(1-\gamma)$ with $\gamma=2$.
Figure 2 shows the result of these computations when $\bar{w}=1$. The function $v_{\hat{\sigma}}$, shown in blue, closely matches the globally optimal value function $v^{*}$ computed via OPI (red dotted line). Similarly, in the second panel, the approximated policy $\hat{\sigma}$ (blue line) closely matches the optimal policy $\sigma^{*}$ (red dotted line). This convergence demonstrates that both the $\hat{\sigma}$-value function $v_{\hat{\sigma}}$ and the policy $\hat{\sigma}$ successfully recover their globally optimal counterparts $v^{*}$ and $\sigma^{*}$, respectively. This is consistent with the result in Theorem 4.2, which states that we can compute a globally optimal policy by solving $\max_{\sigma\in\Sigma}v_{\sigma}(\bar{w})$ at any fixed $\bar{w}\in\mathbbm{R}_{+}$.
In subsequent experiments, we tested the robustness of this outcome to variation in the fixed initial condition $\bar{w}$. We found that, as predicted by theory, the resulting function $v_{\hat{\sigma}}$ again closely approximates $v^{*}$, and the resulting policy $\hat{\sigma}$ again closely approximates $\sigma^{*}$. See Figure 7 in Appendix B for a visualization of one experiment.
In Figure 2, the $y$-axis is the finite time approximation of the infinite horizon lifetime value function $v_{\hat{\sigma}}$ defined in (11) and the $x$-axis is the number of training episodes, which records the number of times the policy $\hat{\sigma}$ is updated according to the $\hat{v}_{\hat{\sigma}}(\bar{w})$ computed from the $N$ Monte Carlo samples. The figure shows consistent improvement in approximated lifetime value from following the policy $\hat{\sigma}$ that improves over the training episodes. The sequence of approximated lifetime value $\hat{v}_{\hat{\sigma}}(\bar{w})$ exhibit a sharp increase in the first half of the training episodes, followed by a stable plateau.  [^2] The sequence of approximated lifetime value converges to a higher value than $v_{\hat{\sigma}}(1)$. This is because the $\hat{v}_{\hat{\sigma}}(\bar{w})$ is a finite time approximation (with rollout $T$) of the infinite horizon lifetime value function $v_{\hat{\sigma}}$. The same phenomenon occurs in Figure 7 and subsequent experiments in the reducible case.
5.2. A Reducible MDP
Now consider a modified version of the optimal saving MDP where both returns and labor income are bounded:
$w^{\prime}=\eta^{\prime}(w-c)+y^{\prime},\quad\eta^{\prime}\in[\underline{\eta},\overline{\eta}],\quad y^{\prime}\in[\underline{y},\overline{y}]$ (13)
with $10<\underline{\eta}<\overline{\eta}<$ 1 and $0\leqslant\underline{y}<\overline{y}<\infty$. For $w\in\mathbbm{R}_{+}$ and a Borel set $B\subseteq\mathbbm{R}_{+}$, the stochastic kernel $P_{\sigma}$ is
$P_{\sigma}(w,B)=\int\mathbbm{1}_{B}(\eta^{\prime}(w-\sigma(w))+y^{\prime})\psi(d\eta^{\prime})\varphi(dy^{\prime})$ (14)
where $\psi$ and $\varphi$ have support on $[\underline{\eta},\overline{\eta}]$ and $[\underline{y},\overline{y}]$ respectively. In this case, we let $\psi$ and $\varphi$ be uniform distributions. For the rest of this section, we always set $[\underline{\eta},\overline{\eta}]=[0.5,0.8]$ and $[\underline{y},\overline{y}]=[1,8]$. At the same time, we deliberately continue to take $\mathbbm{R}_{+}$ as the state space. This leads to a failure of irreducibility, as clarified in the next proposition.
Proposition 5.3.
For any $\sigma\in\Sigma$, the transition kernel $P_{\sigma}$ is neither weakly irreducible nor open set irreducible.
Proof.
Fix $\sigma\in\Sigma$. In view of Lemma 4.3, we only need to show that $P_{\sigma}$ is not weakly irreducible. To this end, let $\pi$ be a distribution on $\mathsf{W}$ that assigns positive probability to open sets. Fix initial wealth $w_{0}\in\mathbbm{R}_{+}$. For any $t$-step transition, let $\alpha_{t}\coloneq w_{t}-\sigma(w_{t})$ be savings at step $t$. Then
$w_{t+1}\leqslant\overline{\eta}\alpha_{t}+\overline{y}\leqslant\overline{\eta}w_{t}+\overline{y}.$
Iterating this inequality $n$ times from $w_{0}$
$w_{n}\leqslant\overline{\eta}^{n}w_{0}+\overline{y}(1+\overline{\eta}+\dots+\overline{\eta}^{n-1})=\overline{\eta}^{n}w_{0}+\overline{y}\frac{1-\overline{\eta}^{n}}{1-\overline{\eta}}.$
Hence, there exists an $M\in\mathbbm{R}_{+}$ such that
$w_{n}<\overline{\eta}w_{0}+\overline{y}\frac{1}{1-\overline{\eta}}<M\quad\forall n\in\mathbbm{N}.$
Let $B=(M,\infty)$. Then $\pi(B)>0$, and $P^{n}_{\sigma}(w_{0},B)=0$ for all $n\in\mathbbm{N}$. This shows that $P_{\sigma}$ is not weakly irreducible. ∎
Figure 3. The upper bound law of motion for wealth
The proof of Proposition 5.2 is partly illustrated in Figure 3. The figure shows the 45 degree line and an upper bound law of motion for wealth, obtained by setting consumption to zero and both shocks to their upper bound. States above the steady state are not reachable from states below the steady state.
Figure 4. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $\bar{w}=1$ vs OPI solutions
Figure 5. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the reducible optimal saving model at $\bar{w}=1$.
For reducible MDPs, Algorithm 1 no longer guarantees convergence to the optimal policy when optimized at a single initial state $\bar{w}\in\mathsf{W}$. This limitation is clearly demonstrated in Figure 5, which shows significant discrepancies between the algorithm’s output and the OPI solution. Figure 5 confirms this limitation, showing that extended training episodes fail to improve the policy’s performance. [^3] The training episode in Figure 5 extends longer than that in Figure 2 because we implement early stopping with a patience parameter of 150 episodes, meaning training terminates after 150 consecutive episodes without improvement in the loss.
Figure 5 provides empirical support for Lemma 4.1. The policy network achieves near-optimal performance at lower wealth levels because these levels are reachable from the initial state under the learned policy $\hat{\sigma}$. However, the performance deteriorates for wealth levels that are not reachable from the initial state given the policy $\hat{\sigma}$. (Reachable range of wealth level in the graph is smaller than in Figure 3 because $\hat{\sigma}(w)>0$ for all $w\in\mathsf{W}$.) We also include the result for $\bar{w}=50$ in Appendix C to show that the same phenomenon related to the reachable subset occurs for larger initial wealth levels.
These results confirm that, when irreducibility fails, the policy $\hat{\sigma}$ computed to maximize local optimality is not guaranteed to be optimal, and the performance of the policy is sensitive to the initial state and the reachable subset of the state space starting from the initial state.
5.3. Firm Entry
The next example provides an extension showing how the results in the paper can be used outside the MDP setting. This is significant because many recent examples of dynamic programs are not MDPs (see, e.g., Sargent and Stachurski (2025a)). The example concerns an optimal stopping problem with state-dependent discounting. We prove that, under the assumptions stated below, local optimality implies global optimality.
Specifically, we consider a stopping problem for a firm choosing if and when to enter a market. Stopping corresponds to entering the market and continuing corresponds to waiting until the next period and then deciding again. If the firm enters at time $t$ it receives one-off profit $\pi(X_{t})$, where $(X_{t})_{t\geqslant 0}$ is a Markov sequence on possibly unbounded interval $\mathsf{X}\subset\mathbbm{R}$ with transition kernel $Q$. If the firm continues then it pays a fixed cost $c$. We admit the possibility that discounting is not constant. (This allows for the fact that firms evaluate future profit opportunities based on cost-of-capital, which fluctuates over time.) Thus, the Bellman equation for the firm is
$v(x)=\max\left\{\pi(x),\;-c+\beta(x)\int v(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\right\},$ (15)
where $\beta(x)\in(0,\infty)$ is the discount factor associated with cost-of-capital in state $x$. (Discount factor realizations above one are admitted in order to accommodate occasionally negative interest rates.)
Let $ib\mathsf{X}$ be the increasing functions in $b\mathsf{X}$ and let $icb\mathsf{X}$ be the continuous functions in $ib\mathsf{X}$. We assume that the discount operator $K$ defined by
$(Kf)(x)=\beta(x)\int f(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\qquad(x\in\mathsf{X},\;f\in b\mathsf{X})$
preserves monotonicity and continuity, in the sense that $K$ is invariant on both $ib\mathsf{X}$ and $icb\mathsf{X}$, and that $r(K)<1$, where $r(K)$ is the spectral radius of $K$. For simplicity, we also assume that $Q$ always has full support on $\mathbbm{R}$, so that $Q(x,B)>0$ for any Borel set $B\subset\mathbbm{R}$ with positive Lebesgue measure. (For example, $(X_{t})$ might be driven by an AR(1) sequence with normal shocks.) Finally, we assume that the profit function $\pi$ is bounded, increasing, and continuous.
A policy is a map $\sigma$ from $\mathsf{X}$ to $\{0,1\}$, with $\sigma(x)=1$ indicating the decision to stop. The lifetime value $v_{\sigma}$ of policy $\sigma$ is the unique fixed point of the operator $T_{\sigma}$ on $b\mathsf{X}$ given by
$(T_{\sigma}\,v)(x)=\sigma(x)\pi(x)+(1-\sigma(x))\left[-c+\beta(x)\int v(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\right].$ (16)
We let $\Sigma$ be the set of all policies $\sigma$ such that $T_{\sigma}$ is invariant on $ib\mathsf{X}$.
As for the MDP case, the value function is denoted $v^{*}$ and defined at each $x\in\mathsf{X}$ by $v^{*}(x)\coloneq\sup_{\sigma\in\Sigma}v_{\sigma}(x)$. A policy $\sigma$ is called optimal if $v_{\sigma}(x)=v^{*}(x)$ for all $x\in\mathsf{X}$. In the proofs below we will make use of the Bellman operator for this problem, which takes the form
$(Tv)(x)=\max\left\{\pi(x),\;-c+\beta(x)\int v(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})\right\}\qquad(x\in\mathsf{X},\;v\in b\mathsf{X}).$
The model has standard optimality properties, which we detail in the theorem below.
Theorem 5.4.
Under the stated assumptions, the following results hold:
  1. (a)
    The value function is the unique solution to the Bellman equation in $b\mathsf{X}$,
  2. (b)
    the value function is increasing and continuous, and
  3. (c)
    at least one optimal policy exists.
We will prove in addition the following result, which holds for any $\sigma\in\Sigma$:
Theorem 5.5.
If there exists an $x\in\mathsf{X}$ such that $v_{\sigma}(x)=v^{*}(x)$, then $\sigma$ is optimal.
We begin the proof of Theorem 5.4 and Theorem 5.5 with a series of lemmas. In what follows, an operator $S$ on $b\mathsf{X}$ is called globally stable when $S$ has a unique fixed point $\bar{v}$ in $b\mathsf{X}$ and $S^{n}v\to\bar{v}$ as $n\to\infty$ for all $v\in b\mathsf{X}$.
Lemma 5.6.
For each $\sigma\in\Sigma$, the operator $T_{\sigma}$ is globally stable on $b\mathsf{X}$.
Proof.
Fix $\sigma\in\Sigma$ and $v,w\in b\mathsf{X}$. Applying the triangle inequality, combined with the fact that $K$ is a positive linear operator, we have $|T_{\sigma}\,v-T_{\sigma}\,w|\leqslant|K(v-w)|\leqslant K|v-w|$. Since $\rho(K)<1$ and $b\mathsf{X}$ is a Banach lattice, by Proposition 4.1 and Theorem 2.1 in Stachurski and Zhang (2021), there exists some $N\in\mathbbm{N}$ such that $T_{\sigma}^{N}$ is a contraction. By Banach fixed point theorem, $T_{\sigma}^{N}$ is globally stable and it follows that $T_{\sigma}$ is globally stable on $b\mathsf{X}$. ∎
The next proof uses several concepts from Sargent and Stachurski (2025b). We refer to that paper for the definitions, which are omitted here for brevity.
Proof of Theorem 5.4.
To prove Theorem 5.4 we apply Theorem 5.5 of Sargent and Stachurski (2025b), which applies to the abstract dynamic program $(b\mathsf{X},\mathbbm{T})\coloneq(b\mathsf{X},\{T_{\sigma}:\sigma\in\Sigma\})$. First, in $b\mathsf{X}$ with its usual partial order, the statement $v_{n}\uparrow v$ is equivalent to $v_{n}(x)\uparrow v(x)$ in $\mathbbm{R}$ for each $x\in\mathsf{X}$. From this fact and the monotone convergence theorem, we obtain $Tv_{n}\uparrow Tv$ whenever $(v_{n})\subset b\mathsf{X}$ and $v_{n}\uparrow v\in b\mathsf{X}$. Thus, $(b\mathsf{X},\mathbbm{T})$ is order continuous. Also, $b\mathsf{X}$ is countably-Dedekind complete because the set of bounded Borel measurable functions is closed under pointwise convergence. In addition, each $T_{\sigma}$ is globally stable (Lemma 5.6) and hence order stable (see Sargent and Stachurski (2025b), Example 2.1). Finally, let $\hat{T}$ be the operator given by $\hat{T}v=|\pi|+Kv$. Since $\rho(K)<1$, there is a $u\in b\mathsf{X}_{+}$ with $\hat{T}u=u$. For this $u$ we have $T_{\sigma}\,u\leqslant\hat{T}u\leqslant u$, so $(b\mathsf{X},\mathbbm{T})$ is bounded above. It now follows from Theorem 5.5 of Sargent and Stachurski (2025b) that the value function is the unique solution to the Bellman equation and at least one optimal policy exists.
To prove the final claim we first note that $T$ is globally stable on $b\mathsf{X}$, since, given $v,w\in b\mathsf{X}$, we have $|Tv-Tw|\leqslant|K(v-w)|\leqslant K|v-w|$, so the argument in the proof of Lemma 5.6 applies. Now fix $v\in icb\mathsf{X}$. Since $K$ is invariant on $icb\mathsf{X}$, the element $-c+Kv$ is in $icb\mathsf{X}$. Moreover, $\pi\in icb\mathsf{X}$ and so $\max\{\pi,-c+Kv\}\in icb\mathsf{X}$. This shows that $T$ is invariant on $icb\mathsf{X}$. Since $icb\mathsf{X}$ is nonempty and closed under uniform limits, and $T$ is globally stable on $b\mathsf{X}$, we conclude that $v^{*}\in icb\mathsf{X}$. The proof of Theorem 5.4 is now done. ∎
Lemma 5.7.
If $v_{\sigma}(x)=v^{*}(x)$ for some $x\in\mathsf{X}$, then $v_{\sigma}=v^{*}$ almost everywhere.
Proof.
Since $v_{\sigma}\leqslant v^{*}$ and $T_{\sigma}\,v\leqslant Tv$ for all $v\in b\mathsf{X}$, we have $v_{\sigma}=T_{\sigma}\,v_{\sigma}\leqslant T_{\sigma}\,v^{*}\leqslant Tv^{*}=v^{*}$. As a result, $v_{\sigma}(x)=v^{*}(x)$ implies $(T_{\sigma}\,v_{\sigma})(x)=(T_{\sigma}\,v^{*})(x)$, which, using the definition of $T_{\sigma}$ in (16), yields
$\beta(x)\int v^{*}(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime})=\beta(x)\int v_{\sigma}(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime}).$
As $\beta(x)>0$, we obtain $\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))Q(x,\mathop{}\!\mathrm{d}x^{\prime})=0$. If $v^{*}>v_{\sigma}$ on a set $E$ of positive Lebesgue measure, then, at the same time, we have
$\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))Q(x,\mathop{}\!\mathrm{d}x^{\prime})\geqslant\int_{E}(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))Q(x,\mathop{}\!\mathrm{d}x^{\prime})>0,$
where the last inequality is due to the assumption that $Q(x,\mathop{}\!\mathrm{d}x^{\prime})$ assigns positive measure to any such $E$. From this contradiction, we conclude that $v^{*}=v_{\sigma}$ almost everywhere. ∎
Lemma 5.8.
Let $f,g$ be two increasing functions in $b\mathsf{X}$ with $f\leqslant g$. If $g$ is continuous and $f=g$ almost everywhere, then $f=g$.
Proof.
If $f$ is continuous, the result follows from the fact that any two continuous measurable functions equal almost everywhere on an interval of $\mathbbm{R}$ are equal. Now suppose instead that $f$ is not continous at $x_{0}\in\mathsf{X}$. Let $a=\lim_{x\uparrow x_{0}}f(x)$ and $b=\lim_{x\downarrow x_{0}}f(x)$. Monotonicity and discontinuiy of $f$ implies $a<b$. Since $g$ is continuous, $g((a,b))^{-1}$ is open. We claim that $g((a,b))^{-1}$ is nonempty. Indeed, since $f=g$ almost everywhere, there are $y,y^{\prime}\in\mathsf{X}$ such that $g(y)\leqslant a$ and $g(y^{\prime})\geqslant b$. Since $g$ is continuous and increasing, there exist $y^{\prime\prime}\in\mathsf{X}$ such that $g(y^{\prime\prime})\in(a,b)$. Thus $g((a,b))^{-1}$ is a nonempty open set. This proves the existence of a set of nonzero Lebesgue measure on which $f<g$. Contradiction. ∎
Proof of Theorem 5.5.
Suppose there exists an $x\in\mathsf{X}$ and a policy $\sigma$ such that $v_{\sigma}(x)=v^{*}(x)$. Then, by Lemma 5.7, we have $v_{\sigma}=v^{*}$ almost everywhere. By Theorem 5.5, the function $v^{*}$ is increasing and continuous. By Lemma 5.6, the operator $T_{\sigma}$ is invariant on the closed set $ib\mathsf{X}$, so the fixed point is in $ib\mathsf{X}$; in particular, $v_{\sigma}$ is also increasing. By definition, $v_{\sigma}\leqslant v^{*}$. Hence, by Lemma 5.8, we have $v_{\sigma}=v^{*}$. ∎
5.4. A Two-State Example
In this section, we show that the irreducibility assumptions used in Theorem 3.3, Theorem 4.2, and Theorem 4.4 cannot be dropped: without irreducibility, (E1)(E3) are not generally equivalent. To make the example as simple as possible, we specialize to the case where $\mathsf{X}$ is finite and study the role of irreducibility in this setting. In applying these theorems, we take the metric on $\mathsf{X}$ to be the discrete metric, so that all subsets of $\mathsf{X}$ are open and $\mathscr{B}$ is all subsets of $\mathsf{X}$. We also assume that $\mathsf{A}$ is finite and impose the discrete metric on $\mathsf{A}$. As a result, every map and hence every policy from $\mathsf{X}$ to $\mathsf{A}$ is continuous.
Prior to stating our example, we note some elementary properties of irreducibility on discrete states. Given a transition kernel $Q$ on $\mathsf{X}$, we say that $y$ in $\mathsf{X}$ is $Q$-accessible from $x\in\mathsf{X}$ when there exists an $m\in\mathbbm{N}$ such that $Q^{m}(x,y)>0$. The usual definition of irreducibility of a transition kernel $Q$ on finite $\mathsf{X}$ is that, for every $x,y\in\mathsf{X}$, $x$ is $Q$-accessible from $y$ and $y$ is $Q$-accessible from $x$. To distinguish between different notions of irreducibility, we call this discrete irreducibility.
Lemma 5.9.
Let $\sigma$ be any feasible policy. When the state space is finite, the following statements are equivalent:
  1. (a)
    $P_{\sigma}$ is strongly irreducible.
  2. (b)
    $P_{\sigma}$ is discretely irreducible.
Proof.
((a) $\implies$ (b)) Fix $x,y\in\mathsf{X}$. Let $\mathbbm{1}_{x}(z)$ equal $1$ when $z=x$ and zero elsewhere. Let $\mathbbm{1}_{y}$ be defined analogously. Since $\mathbbm{1}_{x}\in b\mathsf{X}_{+}$ and $\mathbbm{1}_{y}\in(b\mathsf{X}_{+})^{\prime}$, there exists an $m\in\mathbbm{N}$ with $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle>0$. But $\left\langle{\mathbbm{1}_{x},P_{\sigma}^{m}\mathbbm{1}_{y}}\right\rangle=P_{\sigma}^{m}(x,y)$, so $y$ is $P_{\sigma}$-accessible from $x$. This proves that $P_{\sigma}$ is discretely irreducible.
((b) $\implies$ (a)) Fix a nonzero $f\in b\mathsf{X}_{+}$ and nonzero $\mu\in(b\mathsf{X}_{+})^{\prime}$. Since $\mathsf{X}$ is finite, these element are just maps from $\mathsf{X}$ to $\mathbbm{R}_{+}$ and, as both are nonzero, we can find $\bar{x},\bar{y}\in\mathsf{X}$ such that $f(\bar{y})>0$ and $\mu(\bar{x})>0$. Moreover, since $P_{\sigma}$ is discretely irreducible, there is $m\in\mathbbm{N}$, such that $P_{\sigma}^{m}(\bar{x},\bar{y})>0$. As a result,
$\left\langle{\mu,P_{\sigma}^{m}f}\right\rangle=\sum_{x}(P_{\sigma}^{m}f)(x)\mu(x)=\sum_{x}\sum_{y}f(y)P_{\sigma}^{m}(x,y)\mu(x)\geqslant P_{\sigma}^{m}(\bar{x},\bar{y})f(\bar{y})\mu(\bar{x})>0.$
This proves that $P_{\sigma}$ is strongly irreducible. ∎
Now consider a two-state MDP with $\mathsf{X}=\{1,2\}$ and $\mathsf{A}=\{1,2\}$. The feasible coprrespondence is defined by $\Gamma(1)=\{1,2\}$ and $\Gamma(2)=\{2\}$. The reward function is defined by $r(i,j)=r_{ij}$ with
$\begin{pmatrix}r_{11}&r_{12}\\ r_{21}&r_{22}\end{pmatrix}=\begin{pmatrix}0&1\\ 0&2\end{pmatrix}.$
We set $\beta=0.9$. The transition probabilities $P(x,a,x^{\prime})$ are given by
$P(1,1,\cdot)=(1,0),\quad P(1,2,\cdot)=(1,0),\quad P(2,1,\cdot)=(0,1),\quad P(2,2,\cdot)=(0,1).$
By the definition of the feasible correspondence $\Gamma$, there are only two feasible policies $\Sigma=\{\sigma,\pi\}$, where $\sigma(x)=2$ and $\pi(x)=x$ for all $x\in\mathsf{X}$. The transition probabilities following the two policies are given by
$P_{\sigma}=P_{\pi}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
Observe that $P_{\sigma}=P_{\pi}$ fails to be discretely irreducible. By Lemma 4.3 and Lemma 5.9, this means that this transition kernel is neither strongly irreducible, nor weakly irreducible, nor open set irreducible.
Let us compute the lifetime value functions for the optimal $\sigma$ and $\pi$. For policy $\sigma$, we have $r_{\sigma}=(r_{12},r_{22})=(1,2)$ and $v_{\sigma}=(10,20)$. For $\pi$, we have $r_{\pi}=(r_{11},r_{22})=(0,2)$ and $v_{\pi}=(0,20)$. Since, $v^{*}(x)\coloneq\sup_{s\in\Sigma}v_{s}(x)=v_{\sigma}(x)$, we see that $\sigma$ is an optimal policy. On the other hand, $v_{\pi}(2)=v^{*}(2)$, but $v_{\pi}(1)<v^{*}(1)$, indicating that optimality in one state does not guarantee global optimality. This confirms that irreducibility cannot be dropped from the statements of Theorems 3.3, 4.2, and 4.4.
6. Extensions and Future Work
Using MDP optimality results from Bäuerle and Rieder (2011) or Bertsekas (2022), it should be possible to extend our results to the case of unbounded rewards by replacing the ordinary supremum norm on $b\mathsf{X}$ with a weighted supremum norm. Also, while our results have focused on standard MDPs with constant discount factors, one useful variation of this model is MDPs with state-dependent discount factors, so that $\beta$ becomes a map from $\mathsf{X}$ to $\mathbbm{R}_{+}$ (see, e.g., Stachurski and Zhang (2021)). We conjecture that the main results we obtained for MDPs will extend to generalized ADPs with state-dependent discounting under suitable stability and irreducibility assumptions. Finally, it seems likely that results similar to Theorem 3.3 will be valid for some continuous time MDPs, as well as at least some of the nonstandard discrete time dynamic programs discussed in Bertsekas (2022) and Sargent and Stachurski (2025a). These topics are also left for future work.
Appendix A Algorithm
This section provides a detailed description of the algorithm used to train the policy network described in Section 5.1. In the state transition step, the algorithm uses the wealth bounds $(w_{\min},w_{\max})$ to ensure that the wealth process remains within these bounds to match the OPI solution, which operates on a wealth grid. In practice, we can replace it with $w_{t+1}\leftarrow\eta_{t}(w_{t}-c)+y_{t}$.
We implement the policy network and optimizer using PyTorch. The training is conducted on an NVIDIA RTX 5070 Ti GPU with CUDA 12.8. Our implementation uses Python 3.13.2 and leverages PyTorch’s built-in automatic differentiation (Paszke et al., 2017) to compute gradients of the loss function efficiently.
Input: Discount factor $\beta$, distributions $(\psi,\varphi)$, episodes $K$, rollout steps $T$, batch size $N$, learning rate $\alpha$, wealth bounds $(w_{\min},w_{\max})$
1 Initialize policy network $\hat{\sigma}$ and optimizer
2 for episode $k=1$ to $K$ do
3       Initialize states $w\leftarrow\bar{w}\in\mathbbm{R}^{N}$
4       Pre-sample shocks vectors $\eta_{t}\sim\psi$, $y_{t}\sim\varphi$ for all $t=0,1,\ldots,T-1$
5       for $t=0$ to $T-1$ do
6             $c\leftarrow\hat{\sigma}(w_{t};\theta)$
             $w_{t+1}\leftarrow\min\{\max\{\eta_{t}(w_{t}-c)+y_{t},w_{\min}\},w_{\max}\}$
              // Transitions
7
8       end for
9      $\theta\leftarrow\theta-\alpha\nabla_{\theta}L(\theta)=\theta+\alpha\nabla_{\theta}\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T-1}\beta^{t}u(\hat{\sigma}(w_{i,t};\theta))$
10
11 end for
return $\hat{\sigma}(\cdot;\theta)$
Algorithm 1 Deterministic Policy Gradient
Appendix B Irreducible Optimal Savings MDP with $\bar{w}=50$
Figure 6. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $\bar{w}=50$ against the OPI solutions.
Figure 7. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the irreducible optimal saving model at $\bar{w}=50$.
Appendix C Reducible Optimal Savings MDP with $\bar{w}=50$
In this appendix, we discuss the result of reducible optimal saving MDPs with initial wealth $\bar{w}=50$. As shown in Figure 9, the policy network $\hat{\sigma}$ is able to approximate the optimal policy $\sigma^{*}$ in the reachable subset of the state space. This provides strong empirical evidence for Lemma 4.1.
To emphasize this point, Figure 10 illustrates one sample of the wealth dynamics under the policy network $\hat{\sigma}$ at epoch 750 starting from $\bar{w}=1$ and $\bar{w}=50$. The wealth process initialized at $\bar{w}=50$ explores a substantially broader range of wealth levels compared to the process starting from $\bar{w}=1$. This wider exploration of the state space directly contributes to the improved performance observed in Figure 9.
We note that the two sequences converge to similar patterns after a brief initial period because we used identical random seeds when generating shocks $\eta_{t}$ and $y_{t}$ when training all models. This technical choice minimizes the influence of factors other than the initial wealth level on the comparisons of model performance. The similarity in the wealth trajectories indicates that the policy network gives similar consumption decisions at lower wealth levels.
Nevertheless, the policy still exhibits limitations: since consumption must be strictly positive (i.e., $\hat{\sigma}(w)>0$ for all $w\in\mathsf{W}$), the algorithm struggles to accurately approximate the optimal policy for very high wealth values which are not reachable under $\bar{w}=50$ given the bounded shocks. This discrepancy paired with bounded uniform shocks causes the $\hat{\sigma}$-value function to systematically underestimate the true value function in the upper tail of the wealth state in the left panel of Figure 9.
Figure 8. $\hat{v}_{\sigma}$ and $\hat{\sigma}$ with $\bar{w}=50$ vs OPI solutions
Figure 9. $\hat{v}_{\hat{\sigma}}(\bar{w})$ over training episode for the reducible optimal saving model at $\bar{w}=50$.
These empirical findings emphasize the importance of Lemma 4.1, which formally establishes that when the irreducibility condition fails, the policy network achieves near-optimal performance in reachable subsets of the state space starting from $\bar{w}=50$.
Figure 10. Comparison of wealth trajectories for $\bar{w}=1$ and $\bar{w}=50$ in a typical rollout.